### v1

# Revisiting Global Income Convergence in the 21st Century

Abstract
This paper revisits the debate on income convergence between poor and rich countries. I challenge the view that there is little to no catch-up, and that changes in total factor productivity (TFP) drives cross-country income differences. Since 2000, income levels in poor countries have converged with rich countries at 0.8% annually, rising to 1.5% when excluding Sub-Saharan Africa. A growth accounting exercise incorporating capital income share heterogeneity shows that most convergence since 1980, and over half since 2000 outside Sub-Saharan Africa, results from convergence in physical and human capital inputs rather than TFP.
keywords:
Income Convergence, Capital Accumulation, Growth Accounting
pacs:
JEL Classification]O40, O16, O47
1 Introduction
The growth literature traditionally finds little to no evidence of income levels in poor countries catching up to those in rich countries (unconditional convergence). It further attributes most income disparities across countries to total factor productivity (TFP) differences [12]. Moreover, growth accounting studies often assume a constant capital share of income of one-third across countries. This paper challenges these prevailing views. I find that after 2000, there was unconditional convergence at a rate of 0.8% for a broad sample of countries, and 1.5% when excluding Sub-Saharan Africa. This rate of unconditional convergence is in stark contrast to no unconditional convergence in [3], and higher than the convergence rate of 0.425% for the broad sample of countries and 1.175% outside Sub-Saharan Africa reported in [18]. More importantly, I find that convergence in physical capital and human capital, rather than TFP, has driven all the income convergence since 1980 and more than half of the convergence since 2000 outside Sub-Saharan Africa. Further, I find that capital income shares are not only higher than the commonly assumed one-third but also heterogeneous across countries, following an inverse U-shape across the income distribution.
I find different results compared to the previous literature primarily for two reasons. First, I include the time period post-2000 in my analysis where most of the convergence has actually taken place. In comparison, earlier growth accounting studies were based on a time period where divergence was the norm. Second, I incorporate heterogeneity of capital income shares as observed in the data. I show that assuming a constant capital income share across countries significantly reduces the role of physical capital and human capital in driving convergence, especially for the time period post-2000.
The findings in the paper have important theoretical and policy implications. Economic growth models should incorporate heterogeneity in capital income shares and place greater emphasis on growth driven by the accumulation of capital. Since physical and human capital have been the primary drivers of income convergence, policymakers should prioritize investments in infrastructure, machinery, education, and healthcare, especially in Sub-Saharan Africa where growth has been slow.
This study builds on the papers by [18] and [15], who highlight the shift from divergence to convergence in recent decades. I add to the literature by examining the proximate correlates correlates of convergence within a traditional growth accounting framework, focusing on the roles of physical capital, human capital, and TFP. Earlier growth accounting literature, such as [14], [7], and [12], focused on periods when divergence between countries was the norm, attributing most growth to TFP. By incorporating heterogeneity in capital shares and focusing on a period marked by convergence, I provide new insights into the factors driving income convergence, highlighting the significant roles of physical and human capital accumulation.
2 Data
adjusted for purchasing power parity (PPP) ($rgdpo$), population ($pop$), human capital ($hc$), and labor share of income ($labsh$) for all countries come from the Penn World Tables (PWT10.01). [^2] See [8] for details. Regional classifications are based on the World Bank’s classification of economies. I calculated the capital-output ratio as the ratio of real capital stock (rnna) to real GDP (rgdpna). I use per capita GDP, adjusted to 2017 PPP dollars, as the measure of income ($rgdpo/pop$).
The main sample consists of a balanced panel of 131 countries from 1970 to 2019. The main sample excludes countries with small populations and those whose GDP is heavily dependent on oil rents, which is standard practice in the literature. [^3] Specifically, I exclude countries whose population never exceeded 0.2 million at any point in time. I also exclude countries where oil rents accounted for more than 50% of GDP at any time. The capital measure in the PWT includes fixed reproducible capital but excludes natural resource capital, such as land, inventories, subsoil assets, or intangible capital. This difference is particularly significant for oil-rich countries, which is why they are excluded from the main analysis. [15] also excludes small countries, and oil rich countries in their analysis. I use a sample of 84 economies outside Sub-Saharan Africa, for which the PWT reports human capital, physical capital, and capital share of income to conduct the main growth decomposition exercise. [^4] For variance measures, I further exclude Venezuela. Venezuela is excluded for statistical reasons, as it experienced an economic crisis beginning in 2000, with current GDP per capita lower than its 1960 level. This makes Venezuela an outlier with a large influence on variance measures. The time series of Venezuela’s GDP per capita is plotted in the supplementary appendix for reference.
3 Rate of Income Convergence
This section provides evidence that, post-2000, initially poor countries have been growing faster than initially rich countries. This phenomenon, where poorer economies tend to grow more rapidly than wealthier ones, is commonly referred to as unconditional convergence (or $\beta$-convergence) in the growth literature. I quantify the rate at which poorer countries have been catching up with their richer counterparts. Formally, I calculate the rate of $\beta$-convergence between time period $t$ and $T$ by estimating the following regression:
$\frac{1}{T}(\log(y_{i,T})-\log(y_{i,t}))=\gamma+\beta\log(y_{i,1980})+\epsilon_{i,t}.$ (1)
Here, $y_{i,t}$ refers to the income of country $i$ in year $t$, and $\epsilon_{i,t}$ captures the shocks to the annualized growth rate between time periods $t$ and $T$. I assume that these shocks are independently and identically distributed (iid) with a normal distribution. I estimate separate regressions for the periods 1980-2000 and 2000-2019. In this specification, $\beta$ is identified using cross-sectional variation in annual growth rates across countries initially at different income levels. Additionally, note that the regression in equation 1 includes no country-level controls, meaning that the $\beta$ identified here reflects the rate of unconditional convergence. In both cases, I fix the base-year income at its 1980 level. Supplementary Appendix provides additional analysis, running the convergence regression with 1960 as the base year.
Table 1: Beta Convergence for 1980-2000 and 2000-2019.
All Countries Outside SSA
(1) (2) (3) (4)
1980-2000 2000-2019 1980-2000 2000-2019
$\beta$ 0.005** -0.008*** -0.001 -0.015***
(0.002) (0.002) (0.002) (0.003)
N 131 131 91 91
R2 0.046 0.099 0.001 0.238
Standard errors in parentheses
* $p<0.10$, ** $p<0.05$, *** $p<0.01$
Table 1 presents the estimates for the rate of $\beta-$convergence. Columns (1) and (2) report the estimates for the full balanced panel of countries, while Columns (3) and (4) show results for the set of countries excluding Sub-Saharan Africa. In Column (1), the convergence coefficient is positive and significant for the full sample of countries during the 1980-2000 period. This indicates that rich countries were initially growing faster than poor countries, leading to increased income divergence. However, for countries outside Sub-Saharan Africa, the convergence coefficient is statistically indistinguishable from zero during the same period, providing no evidence of convergence in either sample.
In contrast, the rate of convergence has seen a remarkable reversal in the past two decades. In the full sample, initially poor countries have been catching up to initially rich countries at an annual rate of 0.8%. For countries outside Sub-Saharan Africa, the catch-up rate has been even faster, at 1.5% annually. For comparison, [18] found a the rate of convergence of 0.425% for the full sample of countries and 1.175% outside Sub-Saharan Africa between 2000-2019 with 2000 as the base year. [15] don’t report the rate of convergence outside Sub-Saharan Africa but do find the rate of convergence to be 0.7% in broad sample of countries between 2005-2015 with 2005 as the base year. [1] estimates the rate of conditional convergence to be 1.7% regressing 5-year growth rates on lagged income level and country level-time varying controls. The unconditional convergence rate of 1.5% observed here is substantial in comparison. This rate of $\beta-$convergence implies a half-life of 46 years, meaning that, starting from 2019, poor countries outside Sub-Saharan Africa would close half the income gap with developed countries by 2065, assuming all countries share the same steady-state income level.  [^5] The half-life is calculated as $\tau=-\ln(2)/\ln(1-\beta)$. Moreover, nearly 24% of the variation in growth rates across countries outside Sub-Saharan Africa can be explained by their income levels in 1980. Macro growth regressions typically suffer from low explanatory power since there are a lot of variables determining future growth trajectories not included in the model. The relatively high $R^{2}$ achieved here without the inclusion of additional covariates, suggests that a country’s position in the income distribution is a strong proxy for a range of other factors affecting subsequent growth rates.
Given that most of the convergence post-2000 is observed outside Sub-Saharan Africa, and that data for countries outside Sub-Saharan Africa is generally more reliable [21], the remainder of the paper focuses on this subset of countries. Results for the full set of countries are provided in the Supplementary Appendix.
4 Trends in cross country income differences
This section documents that the decline in overall dispersion of income among counties outside Sub-Saharan Africa started to decline post 2000. Table 2 presents the evolution of income dispersion across countries outside Sub-Saharan Africa over time, using various measures. Columns 1, 2, and 3 display the P90/P10, P90/P50, and P50/P10 percentile ratios of GDP per capita, which capture different aspects of income inequality across the distribution. Column 4 reports the variance of the log of GDP per capita, and Column 5 shows the income ratio between the top 5 and bottom 5 countries.
The data reveal that all measures of income dispersion increased between 1980 and 1990. While the P90/P10 and P50/P10 ratios began to decline after 1990, the most notable trend is the consistent decrease in all measures of income inequality starting from 2000. By 2019, each measure of income dispersion reached its lowest point in the last 40 years.
Table 2: Income Differences across Years (outside SSA)
Year P90/P10 P90/P50 P50/P10 Var(log(GDPpc)) Income Ratio
1980 15.75 3.22 4.89 1.08 35.29
1990 19.81 3.49 5.67 1.17 37.08
2000 17.10 3.87 4.42 1.20 38.30
2010 11.22 2.92 3.85 0.91 34.65
2019 9.97 2.73 3.64 0.81 24.78
Given the significant decline in cross-country income dispersion, the next section conducts an accounting exercise to decompose this convergence into parts attributable to convergence in TFP, the capital-output ratio, and human capital.
5 Accounting for Decline in Income Dispersion
In this section, I perform an accounting exercise to attribute the decline in income dispersion to reductions in TFP differences and input dispersion. Unlike many standard approaches in the literature, which assume a uniform capital share parameter of one-third, this accounting exercise based on percentile ratios allows the capital share of income to vary continuously with the income percentile, as observed in PWT10.01. The accounting exercise using variance as a measure of income dispersion is presented in the supplementary appendix.
Let $p$ denote the income percentile. Suppose that per capita output ($y$) is a function of physical capital ($k$), human capital ($h$), and a residual term ($A$). [^6] It is common to interpret the residual term ($A$) as total factor productivity (TFP), but more broadly, it includes both observable and unobservable factors that affect output beyond the inputs. The output production function, parameterized by $p$, takes the following form:
$y(p)$ $=A(p)(k(p))^{\alpha(p)}(h(p))^{1-\alpha(p)}$
This is the standard Cobb-Douglas production function, with the modification that the capital share ($\alpha$) is allowed to vary across the income distribution. This flexibility is motivated by the large variation in capital share across countries and over time. Figure 1 shows the capital share along the income distribution, as computed from PWT10.01. The capital share is not only heterogeneous but higher than what is typically assumed in the literature. The production function can be rearranged and written in intensive form as follows:
$y(p)=A(p)^{\frac{1}{1-\alpha(p)}}(\frac{k}{y}(p))^{\frac{\alpha(p)}{1-\alpha(p)}}(h(p)).$
Writing the production function in this form reduces the effect of differences in per capita capital ($k$), since per capita capital is allowed to adjust endogenously in response to changes in total factor productivity ($A$). [^7] As argued in [14], under firm optimization, equalization of the marginal product of capital with the rental rate means that an increase in total factor productivity (TFP) leads to a mechanical increase in capital, assuming a constant rental rate. Since my goal is to separately identify the contributions of total factor productivity (TFP) growth and capital accumulation to output growth, I prefer to use the production function written in this form. Given this functional form, the income difference between percentiles $p_{0}$ and $p_{n}$ can be decomposed as follows:
[…]
Here, the second and third terms capture the role of inputs, while the residual term ensures that the accounting identity holds. Appendix A provides a detailed explanation of how each component is computed.
I report the results from the accounting exercise for three percentile ratios-P90/P10, P90/P50, and P50/P10-covering countries outside Sub-Saharan Africa for the years 2000-2019, and 1980-2019. These ratios capture the overall trend in convergence outside Sub-Saharan Africa across the income distribution.
Table 3: Accounting for Income Convergence outside SSA
Percentile Ratio $\Delta_{t}\log(\frac{y(p_{n})}{y(p_{0})})$ Cont. Cont. Cont.
(%) A (%) $\frac{K}{Y}$ (%) $h$ (%)
A: 1980-2000
P90/P10 8 487 -225 -175
P90/P50 18 194 -67 -22
P50/P10 -10 -40 60 90
B: 2000-2019
P90/P10 -54 44 43 11
P90/P50 -34 35 44 26
P50/P10 -20 65 40 -15
C: 1980-2019
P90/P10 -46 -33 89 43
P90/P50 -16 -144 169 81
P50/P10 -30 30 47 20
Table 3 presents the results of the accounting exercise, showing changes in income percentile ratios over three periods: 1980-2000, 2000-2019, and 1980-2019. Positive values in the changes of income percentile ratios signify income divergence, indicating that income inequality between countries has increased, while negative values denote income convergence, meaning that income inequality has decreased. The table breaks down the contributions to these changes from three factors: total factor productivity (TFP) differences, capital-output ratios, and human capital differences. Each row of the table details these contributions across three key percentile ratios-P90/P10, P90/P50, and P50/P10-which represent disparities between the rich (90th percentile) and the poor (10th percentile), the rich and the middle-income group (50th percentile), and the middle-income group and the poor, respectively. The sign of each percentage contribution indicates the direction of its contribution relative to the overall change: a positive percentage means the factor’s change aligns with the total change, while a negative percentage means it opposes the total change. For example, during a period of income divergence (a positive change in the percentile ratio), a positive contribution from TFP differences implies that TFP is also diverged during the same period. Similarly, if the percentile ratio is converging (a negative change), a positive contribution from TFP differences indicates that TFP is also converged during the same period.
Panel A of Table 3 shows that between 1980 and 2000, the P90/P10 and P90/P50 income ratios diverged, while the P50/P10 ratio converged. Although the capital-output ratio and human capital converged between rich and poor economies and between rich and middle-income economies, TFP diverged among these groups, leading to overall income divergence. Notably, the income divergence would have been even greater if inputs had not converged. In contrast, the income difference between middle-income and poor economies converged during this period. The convergence of inputs accounts entirely for the income convergence between middle-income and poor economies, as TFP diverged between these groups during this time.
Panel B reveals that, unlike the 1980-2000 period, income differences converged across the income distribution during 2000-2019. This shift was due to TFP diverging in the earlier period to converging in the later period. More importantly, the reduction in the gaps in the capital-output ratio significantly contributed to the narrowing of income differences during 2000-2019. Specifically, the P90/P10 income ratio declined by 54%, with the decline in the capital-output ratio accounting for 43% of this reduction. Similarly, the P90/P50 ratio declined by 34%, and the P50/P10 ratio declined by 20%. The decline in the capital-output ratio explains 44% of the reduction in P90/P50 differences and 40% of the reduction in P50/P10 differences. These findings highlight the critical role of physical capital in driving convergence during this period. When changes in the capital-output ratio are combined with improvements in human capital, the total contribution of inputs becomes even more pronounced: inputs account for 52% of the reduction in P90/P10 disparities, 65% of the reduction in P90/P50 disparities, and 35% of the reduction in P50/P10 disparities. The overall decline in input dispersion accounts for more than half of the reduction in output disparities between the richest (P90) and poorest (P10) economies, as well as between the richest and middle-income (P50) economies.
Panel C presents the findings for the extended period from 1980 to 2019. During this timeframe, convergence in the capital-output ratio was the primary driver of narrowing income disparities. The P90/P10, P90/P50, and P50/P10 income ratios declined by 46%, 16%, and 30%, respectively. The reduction in the capital-output ratio accounts for 89% of the decrease in P90/P10 income differences, while TFP differences contributed negatively with -33%, indicating divergence. For the P90/P50 ratio, the capital-output ratio explains a striking 169% of the reduction, offsetting the negative contribution from the TFP divergence of -144%. For the P50/P10 ratio, the capital-output ratio accounts for 47% of the convergence, with TFP contributing 30%. When changes in the capital-output ratio are combined with improvements in human capital, inputs account for 132% of the reduction in P90/P10 disparities, 250% of the reduction in P90/P50 disparities, and 77% of the reduction in P50/P10 disparities. These results imply that, had it not been for the significant convergence in the inputs, especially the capital-output ratio, the overall income convergence between the richest and poorer countries would have been substantially lower or potentially nonexistent due to TFP divergence during this period.
5.1 Accounting with constant capital income share
Table 4: Accounting for Income Convergence outside SSA ($\alpha=1/3$)
Percentile Ratio $\Delta_{t}\log(\frac{y(p_{n})}{y(p_{0})})$ Cont. Cont. Cont.
(%) A (%) $\frac{K}{Y}$ (%) $h$ (%)
A: 1980-2000
P90/P10 8 412 -137 -175
P90/P50 18 178 -44 -22
P50/P10 -10 -10 30 90
B: 2000-2019
P90/P10 -54 65 24 11
P90/P50 -34 53 24 26
P50/P10 -20 85 25 -15
C: 1980-2019
P90/P10 -46 4 52 43
P90/P50 -16 -88 100 81
P50/P10 -30 53 27 20
Table 4 presents the main accounting result assuming a constant capital share of one-third. Panel A of Table 4 shows that between 1980 and 2000, the divergence of TFP among rich and poor economies, as well as between poor and middle income economies still accounts for all the divergence in income ratios. This suggests that for studies focused on this time period, such as [14], the assumption of a constant capital share does not make much of a difference. Panel B of Table 4, shows that between 2000 and 2019, TFP accounts for the majority of convergence among rich and poor, middle and poor, and rich and middle income economies. This contrasts with the results in Panel B of Table 3, where inputs accounts for majority of the convergence. Therefore, for the period post-2000, the choice of capital income share becomes crucial. When heterogeneity of capital income shares is incorporated into the growth accounting exercise, capital alone accounts for 43%, 44%, and 40% of the convergence in P90/P10, P90/P50, and P50/P10 ratios between 2000 and 2019. In comparison, with the assumption of a constant capital income share of one-third, the convergence of capital accounts for only 24%, 24%, and 25% of the convergence, respectively. Hence, not incorporating heterogeneity in capital income shares during this time period would lead to the incorrect conclusion that TFP remains the dominant factor driving growth and convergence. Panel C of Table 4 shows that for the extended period from 1980 to 2019, with a constant one-third capital income share, convergence of capital accounts for 52%, and 27% of the convergence in P90/P10 and P50/P10 ratios, respectively, compared to 89% and 47% when heterogeneity of capital income shares is included. Capital still accounts for all the convergence in the P90/P50 ratio in both cases. Therefore, even in the long run, the assumption of a constant capital income share significantly reduces the role of capital in driving income convergence between rich and poor, and between middle-income and poor economies.
Overall, the findings indicate that the convergence of inputs-particularly physical capital-accounts for all the income convergence between rich and poor economies during 1980-2019, and more than half of the convergence during 2000-2019 outside Sub-Saharan Africa. Moreover, the choice of capital income share is important. Assuming a constant capital share of one-third reduces the role of capital in driving convergence by approximately half.
6 Measurement of Inputs and Capital Share of Income
The accounting exercise suggests that the convergence of inputs, particularly physical capital,has significantly driven overall convergence. This section discusses how the measures of physical and human capital are constructed in the Penn World Tables (10.01) and evaluates the extent to which the accounting results depend on the assumptions underlying the construction of these variables.
Physical Capital
In this paper, the capital-output ratio is calculated as the ratio of real capital stock ($rnna$) to real GDP ($rgdpna$) at constant national prices from the Penn World Tables (PWT 10.01).
The capital stock series in the PWT is constructed using the Perpetual Inventory Method (PIM). The stock of capital of type $a$ at time $t$ is computed as the accumulated type-specific investment ($I_{at}$) net of depreciation ($\delta_{a}$):
$K_{at}=I_{at}+(1-\delta_{a})K_{a,t-1}.$
The aggregate capital stock is simply the sum of asset specific capital stocks, $K_{t}=\sum_{a}K_{at}$.
The PIM requires an assumption about the initial capital stock ($K_{0}$), i.e., the capital stock in the first year when investment data becomes available. A common approach in the literature is to use the steady-state relation from the neoclassical growth model:
$K_{0}=\frac{I_{0}}{\delta+g},$
where $g$ is the steady state growth rate of investment. The PWT modifies this approach by using the steady-state assumption initially, then simulating the path of capital for all countries based on the PIM. It calculates cross-country median values of the capital-output ratio ($k_{a}$) and assumes all countries start with the same ratio. The initial capital stock is then recalculated as:
$K_{a0}=Y_{0}k_{a0}.$
This implies that differences in initial capital stocks are solely due to differences in initial output levels. However, this assumption is unlikely to affect the accounting results significantly. For countries in the main analysis, investment data is available since 1970 or earlier. Over time, the influence of the initial capital stock diminishes due to depreciation, especially for assets with high depreciation rates. Assuming a 5% annual depreciation rate, on average only 5.7% of capital in 2000 is due to underappreciated capital in 1970 for the countries included in the main analysis. [^8] The capital in year $1970+k$ is given by $K_{1970+k}=\sum_{t=1}^{t=k}(1-\delta)^{t-1}I_{t}+(1-\delta)^{k}K_{1970}$. The amount of capital in year $1970+k$ due to undepreciated capital in year $1970$ is therefore given by $\frac{(1-\delta)^{k}K_{1970}}{K_{1970+k}}$. The amount of capital in year 1980, 1990, 2000, 2010, due to undepreciated capital in 1970 is 32.7%, 13.6%, 5.7%, 2.5% on average.
Therefore, the decline in capital-output ratio differences as a key factor in reducing output differences post-2000 is not driven by assumptions about the initial capital stock. Instead, it is more likely driven by changes in investment patterns rather than initial capital stock differences.
Human Capital
The human capital measure in PWT10.01 is constructed using data on average years of schooling in the population aged 25 and older from [2], [5], and assumed rates of return for primary, secondary, and tertiary education based on Mincer equation estimates from [19] and [4]. Specifically:
$hc=exp\{\phi(s)\}$
where
$\phi(s)=\begin{cases}0.134\cdot s&\text{if }s\leq 4;\\ 0.134\cdot 4+0.101\cdot(s-4)&\text{if }4<s\leq 8;\\ 0.134\cdot 4+0.101\cdot 4+0.068\cdot(s-8)&\text{if }s\geq 8\end{cases}$
This measure of human capital is imperfect because it does not account for the quality of schooling [20, 10] or cross-country differences in returns to experience [16]. Since the returns to schooling are assumed to be constant across years, human capital convergence is solely driven by the convergence in years of schooling. If returns to schooling, the quality of education, or returns to experience are also converging across economies, then the contribution of human capital presented here is a lower bound.
Measurement of Capital Share of Income
The capital share of income ($\alpha$) in PWT10.01 is derived using the following relationship:
$\alpha=1-labsh$
where labsh represents the share of labor income in GDP. One significant challenge in measuring the labor share is that the labor income of self-employed workers is not directly observable. [9] discusses various methods for estimating the labor compensation of self-employed workers. Two common adjustments rely on mixed-income data from national accounts. Mixed income refers to the total income earned by self-employed workers, encompassing both labor and capital income. The first adjustment allocates all mixed income to labor, while the second adjustment allocates mixed income between labor and capital in the same proportion as the rest of the economy. In cases where mixed income data is unavailable, a third adjustment uses data on the number of employees and self-employed workers, assuming that the self-employed earn the same average wage as employees. PWT10.01 introduces a fourth adjustment that adds the value-added from agriculture to labor compensation of employees. Whenever mixed income data is available, PWT10.01 applies the second adjustment. In the absence of mixed-income data, I use the minimum of the third and fourth adjustments.
Figure 1: Variation of alpha over income percentile.
Figure 1 illustrates the capital share of income across countries at different percentiles of the income distribution for various years. For the sub-sample of countries included in this analysis, PWT shows an average capital share of income of 0.46, which significantly exceeds the commonly assumed value of 1/3 in the literature. Moreover, the capital share of income follows an inverse U-shaped pattern across the income distribution, with middle-income countries having the highest capital income share, followed by poor countries, and high-income countries having the lowest capital income share. Only the richest countries have capital income shares close to 1/3, suggesting that using a constant capital income share may not be appropriate for countries at other points in the income distribution. The decomposition approach in the current paper accommodates this heterogeneity, allowing the capital income share to vary along the income distribution. Additionally,, over time, countries across the income distribution have experienced an increase in the capital share of income, which indicates a corresponding decline in labor share. The recent work by [13] examines this global decline in labor share and attributes much of the trend to cheaper information and communication technology (ICT) capital. The growing importance of the capital share of income reinforces the hypothesis that the convergence of capital has played a substantial role in the convergence of output post-2000.
7 Conclusion
This study revisited the dynamics of cross-country income convergence using the most recent data from the Penn World Tables (PWT 10.01). Roughly 70% of cross-country empirical work relies on the Penn World Tables [11]. The updated PWT data no longer support some of the basic growth accounting facts that underpinned earlier studies. Contrary to the traditional view that emphasizes non-convergence and the limited role of factor inputs, the analysis reveals a significant shift in global income patterns since 2000. There has been unconditional convergence, particularly outside Sub-Saharan Africa. More importantly, accumulation of inputs, particularly physical capital, has significantly driven income convergence outside Sub-Saharan Africa.
This suggests that the factors determining global income dynamics have changed dramatically, necessitating a reevaluation of existing growth theories and models. The results open several avenues for future research. What factors have contributed to the reduction in input dispersion? What underlies the persistent stagnation of Sub-Saharan economies despite their market-oriented reforms? The findings also have significant policy implications. Since input accumulation, rather than TFP growth, primarily drives convergence, policies aimed at removing barriers to the accumulation of physical and human capital in poorer economies may yield more substantial growth dividends than those focused solely on TFP enhancements.
Appendix A Mathematical Derivations
A.1 Generalized Decomposition
Let $p$ denote percentile and output be some arbitrary function if inputs, TFP parametrized by $p$.
$y(p)=F(A(p),k/y(p),h(p)).$
Then,
[…]
$\log(y(p_{n}))-\log(y(p_{0}))$ $=\int_{p_{0}}^{p_{n}}[\frac{d\log F(.)}{d\log A(p)}\frac{d\log A(p)}{dp}+\frac{d\log F(.)}{d\log k/y(p)}\frac{d\log k/y(p)}{dp}+\frac{d\log F(.)}{d\log h(p)}\frac{d\log h(p)}{dp}]$
[…]
where
$\mathcal{E}_{A}(p)$ $=\frac{d\log F}{d\log A}\frac{d\log A}{dp}$
$\mathcal{E}_{k/y}(p)$ $=\frac{d\log F}{d\log(k/y)}\frac{d\log(k/y)}{dp}$
$\mathcal{E}_{h}(p)$ $=\frac{d\log F}{d\log(h)}\frac{d\log(h)}{dp}$
Now suppose that the output production has the following functional form:
$y(p)$ $=A(p)(k(p))^{\alpha(p)}(h(p))^{1-\alpha(p)}$
$=A(p)^{\frac{1}{1-\alpha(p)}}(k/y(p))^{\frac{\alpha(p)}{1-\alpha(p)}}h(p)$
In this case:
$\mathcal{E}_{A}(p)$ $=\frac{1}{1-\alpha(p)}\frac{d\log A}{dp}$
$\mathcal{E}_{k/y}(p)$ $=\frac{\alpha(p)}{1-\alpha(p)}\frac{d\log(k/y)}{dp}$
$\mathcal{E}_{h}(p)$ $=\frac{d\log(h)}{dp}$
The contribution of $A$ can be calculated as a residual. The contribution of each of the inputs can then be evaluated as follows:
[…]
$\approx\sum_{k=0}^{n-1}\frac{\frac{\alpha(p_{k+1})}{1-\alpha(p_{k+1})}+\frac{\alpha(p_{k})}{1-\alpha(p_{k})}}{2}[\log(k/y)(p_{k+1})-\log(k/y)(p_{k})]$
[…]
$=\log(h(p_{n}))-\log(h(p_{0}))$
[…]
Note that if the elasticity parameter is constant along income distribution ( $\alpha(p)=\alpha$), the the generalized decomposition reduces to the usual Hall-Jones decomposition.
$\ln(y(p_{n}))-\ln(y(p_{0}))$ $=\frac{1}{1-\alpha}\ln(\frac{A(p_{n})}{A(p_{0})})+\frac{\alpha}{1-\alpha}\ln(\frac{k/y(p_{n})}{k/y(p_{0})})+\ln(\frac{h(p_{n})}{h(p_{0})}).$
Appendix B Robustness and Additional Exercises
B.1 Unconditional Convergence with 1960 Base Year
Table 5: Beta Convergence for 1980-2000 and 2000-2019.
All Countries Outside SSA
(1) (2) (3) (4)
1980-2000 2000-2019 1980-2000 2000-2019
$\beta$ 0.002 --0.009∗∗∗ --0.004 --0.014∗∗∗
(0.003) (0.003) (0.002) (0.004)
N 107 107 72 72
R2 0.007 0.090 0.034 0.157
Standard errors in parentheses
* $p<0.10$, ** $p<0.05$, *** $p<0.01$
Table 5 presents the results of $\beta$-convergence regressions, using 1960 as the base year for initial income levels. The key objective here is to determine whether the choice of base year affects the conclusions regarding income convergence.
The results confirm that the findings are robust to the use of a different base year. For both the full sample of countries and the sub-sample excluding Sub-Saharan Africa (SSA), there is no evidence of convergence between 1980 and 2000. In this period, the $\beta$ coefficient is close to zero and statistically insignificant in both samples, indicating no relationship between initial income levels and subsequent growth rates.
In contrast, the period 2000-2019 shows strong evidence of convergence, with statistically significant negative $\beta$ coefficients for both the full sample and the sub-sample excluding SSA. The magnitude of convergence is similar to that found when using 1980 as the base year. For the full sample of countries, the $\beta$ coefficient is --0.009, while for countries outside SSA, it is --0.014. These results suggest that, post-2000, poorer countries have been catching up with richer countries at similar rates, regardless of whether 1960 or 1980 is used as the starting point for income levels.
B.2 Robustness to different measures of GDP
Table 6: Beta Convergence for 2000-2019 (Excluding SSA)
RGDPO RGDPE CGDPO CGDPE
$\beta$-0.015∗∗∗ --0.014∗∗∗ --0.015∗∗∗ --0.014∗∗∗
(0.003) (0.003) (0.003) (0.003)
N 91 91 91 91
R2 0.238 0.221 0.250 0.221
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
The Penn World Table (PWT) provides multiple measures of GDP, converted into common currency units, using different approaches to national income accounting. Two measures of real GDP are reported: one based on the output side (RGDPO) and another based on the expenditure side (RGDPE). Similarly, two measures of current-price GDP are available: the output-side measure (CGDPO) and the expenditure-side measure (CGDPE).
Table 6 presents a robustness check of the $\beta$-convergence results for the period 2000-2019, using different measures of per capita GDP for countries outside Sub-Saharan Africa. Each column corresponds to a different measure of GDP used in the calculation of per capita income.
The results indicate that the rate of convergence remains remarkably stable across all four GDP measures. In each case, the $\beta$ coefficient is negative and statistically significant at the 1% level, with convergence rates ranging between --0.014 and --0.015. This consistency suggests that the pattern of income convergence for countries outside Sub-Saharan Africa is robust to the choice of GDP measure, whether based on real output, real expenditure, or current-price accounting. The $R^{2}$ values also indicate a similar explanatory power across all specifications, with values ranging from 0.221 to 0.250.
B.3 Income difference for all countries
Table 7: Income Differences across Years (Full Sample)
Year P90/P10 P90/P50 P50/P10 Var(log(GDPpc)) Income Ratio
1980 18.56 5.30 3.50 1.15 18.49
1990 24.95 5.55 4.50 1.37 23.26
2000 31.95 6.88 4.64 1.62 29.53
2010 27.63 4.51 6.12 1.54 30.71
2019 25.06 4.41 5.69 1.47 27.17
[…]
Table 7 presents income differences across various measures for the full sample of countries, complementing the analysis of countries outside Sub-Saharan Africa (SSA) provided in the main text. The table reports several key indicators of income dispersion, including the P90/P10, P90/P50, and P50/P10 percentile ratios of GDP per capita, the variance of the log of GDP per capita, and the income ratio between the top 5 and bottom 5 of countries.
A key observation from this table is that the decline in income disparities post-2000 has been relatively modest for the full sample of countries, particularly compared to the more substantial reductions observed in countries outside SSA. For instance, the P90/P10 ratio fell from 31.95 in 2000 to 25.06 in 2019, indicating some reduction in inequality between the richest and poorest economies. However, this decline is less pronounced compared to the convergence seen outside SSA, as shown in the main text.
Notably, the P50/P10 ratio increased from 4.64 in 2000 to 5.69 in 2019, reflecting a widening income gap between middle-income and poor economies. This increase is largely driven by the stagnation of many Sub-Saharan economies, which have struggled to keep pace with global growth. Consequently, while income disparities have narrowed somewhat at the top end of the distribution (as seen in the decline in the P90/P10 and P90/P50 ratios), the lower half of the income distribution has experienced greater divergence, especially between middle-income and poorer countries.
B.4 Accounting for Income Convergence (Full Sample)
Table 8: Accounting for Income Convergence (Full Sample)
Year $\Delta_{t}\log(y(p_{n})/y(p_{0}))$ Contribution Contribution Contribution
(%) A (%) K/Y (%) $h$ (%)
Panel A: 1980-2000
P90/P10 54 131 -24 -6
P90/P50 15 247 -67 -87
P50/P10 39 85 -10 23
Panel B: 2000-2019
P90/P10 -27 81 0 19
P90/P50 -29 48 24 24
P50/P10 3 -233 267 67
Panel C: 1980-2019
P90/P10 27 181 -48 -30
P90/P50 -14 -164 121 143
P50/P10 42 62 10 26
Table 8 presents the accounting decomposition for income differences across the global income distribution for three periods: 1980-2000, 2000-2019, and 1980-2019. The analysis includes all economies, including Sub-Saharan Africa.
Panel A of Table 8 shows that between 1980 and 2000, income disparities widened across the income distribution. The P90/P10 income ratio increased by 54%, indicating divergence between the richest and poorest countries. TFP differences contributed significantly to this divergence, accounting for 131% of the increase. Convergence in the capital-output ratio and human capital partially offset this effect, contributing -24% and -6%, respectively. Similarly, the P90/P50 ratio increased by 15%, with TFP differences contributing 247% to divergence, while capital-output ratios and human capital contributed negatively, offsetting some of the divergence. The P50/P10 ratio increased by 39%, mainly due to TFP differences (85%), with minimal offset from inputs.
Panel B of Table 8 reveals that between 2000 and 2019, income disparities decreased between the richest and poorest countries. The P90/P10 income ratio declined by 27%, indicating convergence. TFP differences contributed 81% to this convergence, while human capital contributed 19%. The capital-output ratio had a negligible effect in this case. The P90/P50 ratio decreased by 29%, with TFP and inputs contributing equally to convergence. Notably, the P50/P10 ratio showed a slight divergence of 3%. This divergence was due to a negative contribution from TFP (-233%), but substantial convergence in the capital-output ratio (267%) and human capital (67%) more than compensated for this, indicating significant input-driven convergence among middle-income and poor countries.
Panel C of Table 8 presents the findings for the entire period from 1980 to 2019. Over this timeframe, the P90/P10 income ratio increased by 27%, indicating overall divergence between the richest and poorest countries. TFP differences were the main driver, contributing 181% to this divergence. Convergence in the capital-output ratio (-48%) and human capital (-30%) partially mitigated the divergence caused by TFP. The P90/P50 ratio decreased by 14%, reflecting convergence between the richest and middle-income countries. This convergence was driven by improvements in the capital-output ratio (121%) and human capital (143%), which offset the divergence from TFP differences (-164%). The P50/P10 ratio increased by 42%, indicating divergence between middle-income and poor countries, primarily due to TFP differences (62%).
B.5 Variance Decomposition
This section complements the main accounting analysis by performing the accounting based on variance as a measure of income dispersion, which is more commonly used in the literature. However, I prefer the percentile decomposition over variance decomposition for several reasons. First, percentile ratios allow for continuous variation in parameters across time and countries. Second, percentile ratios are less sensitive to outliers compared to variance measures. Furthermore, variance decomposition includes a term for the covariance between total factor productivity (TFP) and inputs, which can be difficult to interpret. For this exercise, I use a Cobb-Douglas production function, and a capital share parameter ($\alpha$) of 0.46, which is the observed mean in PWT10.01:
Suppose we have a Cobb-Douglas production function.
$Y=\tilde{A}K^{\alpha}(hL)^{1-\alpha}$
This can be rewritten as
$\frac{Y}{L}=A(\frac{K}{Y})^{\frac{\alpha}{1-\alpha}}(h)$
$y=Ay^{kh}$
where $y=\frac{Y}{L}$, $y^{kh}=\frac{K}{Y})^{\frac{\alpha}{1-\alpha}}(h)$. Consequently,
$Var(\ln(y))=Var(\ln(A))+Var(\ln(y^{kh}))+2Cov(\ln(A),\ln(y^{kh}))$
Table 9: Accounting for Variance Decline
Years $\Delta_{t}Var(\ln(y))$ $\Delta_{t}Var(\ln(A))$ $\Delta_{t}Var(\ln(y^{kh}))$ $\Delta_{t}$ 2$\times$×Cov(A,$y^{kh}$)
Outside SSA
1980-2000 11 141 -126 86
2000-2019 -39 27 36 36
1980-2019 -27 -20 104 16
Full Sample
1980-2000 51 69 -20 51
2000-2019 -19 66 43 -9
1980-2019 31 71 -59 88
Table 9 presents the variance decomposition of log income per worker ($\ln(y)$) for both countries outside Sub-Saharan Africa (SSA) and the full sample over three periods: 1980-2000, 2000-2019, and 1980-2019.
For countries outside SSA, between 1980 and 2000, the variance of log income per worker increased by 11%, indicating a slight divergence in income levels among these countries. This increase was primarily driven by a substantial rise in the variance of log TFP, which surged by 141%, suggesting that differences in total factor productivity widened significantly during this period. The covariance term also increased by 86%, further contributing to divergence. However, the variance of log inputs decreased by 126%, indicating strong convergence in physical and human capital accumulation. Had it not been for the convergence in inputs, income variance would have been much higher, as the divergence from TFP differences and the covariance effect would have led to an even greater increase in income disparities.
Between 2000 and 2019, the variance of log income per worker outside SSA decreased by 39%, reflecting a significant convergence in income levels. Decline in variance of inputs accounted for 36% of convergence. The contribution of inputs increases to 54% when including half the covariance term.
The role of inputs is even more pronounced for the time period 1980-2019. During this period the variance of income decline by 27%, with decline in variance of inputs accounting for 104% of the decline, while TFP diverged during this period.
This shows that inputs account for almost half of decline since 2000, and almost all of decline since 1980 in variance of income outside SSA.
For the full sample of countries, between 1980 and 2000, the variance of log income per worker increased by 51%. This increase was primarily driven by a rise in the variance of log TFP by 69%. The covariance term also increased by 51%, further contributing to divergence. However, the variance of log inputs decreased by 20%. Between 2000 and 2019, the variance of log income per worker in the full sample decreased by 19%. Reduction in variance of TFP contributed to 66% of the decline, while reduction in variance of inputs contributed to 43% of decline. Covariance term increased by 9%, contributing to divergence. Over the entire period from 1980 to 2019, the variance of log income per worker in the full sample increased by 31%, indicating a net divergence in global income levels over the long term. This divergence was primarily driven by a substantial rise in the variance of log TFP, which increased by 71%, and an increase in the covariance term by 88%, both significantly contributing to divergence. However, the variance of log inputs decreased by 59%, indicating strong convergence in physical and human capital accumulation.
B.6 Capital Accumulation in Sub-Saharan Africa
It is puzzling to observe a persistent lack of convergence in Sub-Saharan African economies. Despite substantial improvements in policy outcomes across the region, as documented by [6], this has not translated into faster economic growth compared to economies outside Sub-Saharan Africa. [6] measures poor policy outcomes using indicators such as high inflation, significant black market premiums, and negative interest rates, among others. While these measures have improved considerably, growth in Sub-Saharan Africa remains sluggish relative to other regions. Although a comprehensive investigation into the reasons behind this lack of convergence is beyond the scope of the current study, I find that the slow rate of input accumulation is likely a key contributing factor. [17] further highlight capital flight from Sub-Saharan economies to the rest of the world, driven by economic and political uncertainties, which may exacerbate the region’s stagnation.
Figure 2: Capital-Output Ratio across regions
Figure 2 illustrates the mean capital-output ratio, weighted by population size, across different regions. While the capital-output ratio has continued to rise in regions outside Sub-Saharan Africa, Sub-Saharan Africa is the only region where a decline in the capital-output ratio is observed after 2000. This decline likely reflects the region’s challenges in accumulating capital, which may be linked to political instability, economic uncertainty, and capital flight. The capital-output ratio in Figure 2 is computed using a sample of countries for which investment data has been available since 1970 or earlier, ensuring consistent comparisons across regions.
B.7 Influence of Venezuela
Figure 3: Log GDP per capita (2017\$ PPP) for Venezuala
Figure 3 illustrates the significant decline in Venezuela’s GDP per capita, measured in 2017 PPP dollars, following 2010. The sharp drop in Venezuela’s GDP per capita is a direct consequence of the country’s political and economic crisis. Given the unique circumstances surrounding Venezuela’s economic decline, it stands as an outlier in cross-country analyses. Therefore, I exclude Venezuela from the calculation of income variance, as variance is particularly sensitive to outliers.

### v2

@@
 Abstract
-This paper revisits the debate on income convergence between poor and rich countries.
-I challenge the view that there is little to no catch-up, and that changes in total factor productivity (TFP) drives cross-country income differences.
-Since 2000, income levels in poor countries have converged with rich countries at 0.8% annually, rising to 1.5% when excluding Sub-Saharan Africa.
-A growth accounting exercise incorporating capital income share heterogeneity shows that most convergence since 1980, and over half since 2000 outside Sub-Saharan Africa, results from convergence in physical and human capital inputs rather than TFP.
+Recent research has documented a reversal from divergence to convergence in income levels between rich and poor countries after 2000.
+This paper employs a growth accounting framework to investigate the proximate sources of convergence over the period 1980-2019.
+I find that while output levels began to converge only after 2000, capital (physical and human) had already been converging during 1980-2000.
+The divergence in total factor productivity (TFP) during this earlier period offset the gains from capital convergence, resulting in little overall income convergence.
+After 2000, capital maintained its pattern of convergence, and, unlike in the earlier period, TFP also began to converge.
+Quantitatively, more than half of the convergence between 2000 and 2019—and nearly all the convergence between 1980 and 2019 outside Sub-Saharan Africa—can be attributed to the convergence in capital rather than to improvements in TFP.
+These findings highlight that factor accumulation has played a much larger role in driving long-run convergence dynamics than previously documented in the literature.
 keywords:
-Income Convergence, Capital Accumulation, Growth Accounting
+Income Convergence, Capital Accumulation, Growth Accounting, Total Factor Productivity
 pacs:
@@
 1 Introduction
-The growth literature traditionally finds little to no evidence of income levels in poor countries catching up to those in rich countries (unconditional convergence).
-It further attributes most income disparities across countries to total factor productivity (TFP) differences [12].
-Moreover, growth accounting studies often assume a constant capital share of income of one-third across countries.
-This paper challenges these prevailing views.
-I find that after 2000, there was unconditional convergence at a rate of 0.8% for a broad sample of countries, and 1.5% when excluding Sub-Saharan Africa.
-This rate of unconditional convergence is in stark contrast to no unconditional convergence in [3], and higher than the convergence rate of 0.425% for the broad sample of countries and 1.175% outside Sub-Saharan Africa reported in [18].
-More importantly, I find that convergence in physical capital and human capital, rather than TFP, has driven all the income convergence since 1980 and more than half of the convergence since 2000 outside Sub-Saharan Africa.
-Further, I find that capital income shares are not only higher than the commonly assumed one-third but also heterogeneous across countries, following an inverse U-shape across the income distribution.
-I find different results compared to the previous literature primarily for two reasons.
-First, I include the time period post-2000 in my analysis where most of the convergence has actually taken place.
-In comparison, earlier growth accounting studies were based on a time period where divergence was the norm.
-Second, I incorporate heterogeneity of capital income shares as observed in the data.
-I show that assuming a constant capital income share across countries significantly reduces the role of physical capital and human capital in driving convergence, especially for the time period post-2000.
-The findings in the paper have important theoretical and policy implications.
-Economic growth models should incorporate heterogeneity in capital income shares and place greater emphasis on growth driven by the accumulation of capital.
-Since physical and human capital have been the primary drivers of income convergence, policymakers should prioritize investments in infrastructure, machinery, education, and healthcare, especially in Sub-Saharan Africa where growth has been slow.
-This study builds on the papers by [18] and [15], who highlight the shift from divergence to convergence in recent decades.
-I add to the literature by examining the proximate correlates correlates of convergence within a traditional growth accounting framework, focusing on the roles of physical capital, human capital, and TFP.
-Earlier growth accounting literature, such as [14], [7], and [12], focused on periods when divergence between countries was the norm, attributing most growth to TFP.
-By incorporating heterogeneity in capital shares and focusing on a period marked by convergence, I provide new insights into the factors driving income convergence, highlighting the significant roles of physical and human capital accumulation.
+The traditional literature on economic growth has long debated whether poorer countries’ incomes converge toward those of richer countries.
+Empirical studies, particularly those prior to the early 2000s, generally found limited or no support for unconditional convergence [3, 18].
+However, recent studies by [17] and [14] challenge this view, documenting a reversal from divergence to convergence in recent decades.
+This documented reversal prompts two critical yet relatively unexplored questions: first, what factors can explain this turnaround in convergence dynamics; and second, which proximate determinants have contributed significantly to the observed income convergence?
+I address these questions by first replicating the empirical finding of convergence in recent decades, thereby establishing a consistent empirical basis for further analysis.
+I then examine the proximate drivers of this shift using a growth accounting framework that decomposes income growth into contributions from total factor productivity (TFP), physical capital, and human capital.
+In doing so, I explicitly account for heterogeneity in capital income shares across countries, a feature that standard growth accounting exercises often abstract from.
+This approach allows for a more accurate quantification of the role played by each factor in driving convergence, and highlights how their relative contributions have changed over time.
+I find that the reversal from divergence to convergence can be attributed to two main factors: first, total factor productivity (TFP) shifted from a diverging pattern before 2000 to convergence thereafter; and second, physical and human capital inputs continued to converge consistently.
+While TFP transitioned from divergence to convergence post-2000, its contribution to long-run income convergence remains secondary.
+I show that over the period 2000-2019, more than half of the observed income convergence among countries outside Sub-Saharan Africa (SSA) arises from convergence in physical and human capital.
+More strikingly, over the longer timeframe of 1980-2019, income convergence among countries outside SSA occurs despite divergence in total factor productivity (TFP).
+The entirety of the long-run convergence in output levels for these countries is driven solely by narrowing gaps in physical and human capital accumulation.
+Additionally, I find that for poorer countries within Sub-Saharan Africa (SSA), there has been no convergence in the capital-output ratio since 2000.
+Notably, while the extent of total factor productivity (TFP) convergence between rich economies and poor countries outside SSA mirrors that observed between rich economies and poor countries in SSA, the markedly lower degree of output convergence in the latter is entirely accounted for by the absence of capital convergence.
+This pattern suggests that the lack of capital convergence is an important factor contributing to the region’s relative economic stagnation.
+These findings provide a revised understanding of the drivers behind global income dynamics.
+Prior growth accounting and development accounting exercises consistently highlighted differences in total factor productivity (TFP) as the central factor explaining persistent income gaps and divergence across countries [13, 6, 11].
+While TFP indeed remains the dominant explanatory factor in the divergence era preceding 2000, its relative importance diminishes significantly thereafter.
+Specifically, for the period following 2000, and even more clearly over the longer span from 1980 to 2019, the observed convergence in global income levels is primarily driven by convergence in physical and human capital accumulation rather than TFP.
+The differences between these findings and the prior literature primarily reflect two key factors: first, this paper explicitly analyzes the more recent period of convergence post-2000, whereas earlier studies predominantly considered periods of divergence; second, the analysis incorporates heterogeneity in capital income shares across countries and over time.
+Assuming a constant capital income share underestimates the contribution of factor inputs, especially in the period following 2000.
+This paper also contributes to the literature on economic growth in Sub-Saharan Africa (SSA) by identifying the non-convergence of capital accumulation as a critical factor underlying the region’s persistent income stagnation.
+The findings in this paper have important theoretical and policy implications.
+Economic growth models should incorporate heterogeneity in capital income shares and place greater emphasis on growth driven by the accumulation of physical and human capital.
+Given that these factors have been the primary drivers of income convergence, policymakers should prioritize investments in infrastructure, machinery, education, and healthcare—especially in regions such as Sub-Saharan Africa where growth has been slow.
 2 Data
-adjusted for purchasing power parity (PPP) ($rgdpo$), population ($pop$), human capital ($hc$), and labor share of income ($labsh$) for all countries come from the Penn World Tables (PWT10.01).
+This section summarizes the data used in the current study.
+Data for real GDP adjusted for purchasing power parity (PPP) ($rgdpo$), population ($pop$), human capital ($hc$), and labor share of income ($labsh$) for all countries come from the Penn World Tables (PWT10.01).
 [^1] See [7] for details.
 Regional classifications are based on the World Bank’s classification of economies.
-I calculated the capital-output ratio as the ratio of real capital stock (rnna) to real GDP (rgdpna).
+I calculate the capital-output ratio as the ratio of real capital stock (rnna) to real GDP (rgdpna).
 I use per capita GDP, adjusted to 2017 PPP dollars, as the measure of income ($rgdpo/pop$).
-The main sample consists of a balanced panel of 131 countries from 1970 to 2019.
-The main sample excludes countries with small populations and those whose GDP is heavily dependent on oil rents, which is standard practice in the literature.
+The selection of output side real gdp (rgdpo) over alternative Penn World Table output measures follows the recommendations of [7] for comparing productive capacities across countries and over time.
+All samples excludes countries with small populations and those whose GDP is heavily dependent on oil rents, which is standard practice in the literature.
 [^2] Specifically, I exclude countries whose population never exceeded 0.2 million at any point in time.
@@
 The time series of Venezuela’s GDP per capita is plotted in the supplementary appendix for reference.
-3 Rate of Income Convergence
-This section provides evidence that, post-2000, initially poor countries have been growing faster than initially rich countries.
-This phenomenon, where poorer economies tend to grow more rapidly than wealthier ones, is commonly referred to as unconditional convergence (or $\beta$-convergence) in the growth literature.
-I quantify the rate at which poorer countries have been catching up with their richer counterparts.
-Formally, I calculate the rate of $\beta$-convergence between time period $t$ and $T$ by estimating the following regression:
-$\frac{1}{T}(\log(y_{i,T})-\log(y_{i,t}))=\gamma+\beta\log(y_{i,1980})+\epsilon_{i,t}.$ (1)
-Here, $y_{i,t}$ refers to the income of country $i$ in year $t$, and $\epsilon_{i,t}$ captures the shocks to the annualized growth rate between time periods $t$ and $T$.
+3 Income Convergence
+I replicate the empirical framework of [17] and [14] to investigate whether, in the post-2000 period, economies with lower initial income per capita have tended to grow more rapidly than those with higher initial incomes.
+This pattern-commonly referred to as unconditional (or $\beta$-) convergence-implies that poorer countries catch up to richer ones over time.
+To quantify the speed of convergence between dates $t$ and $T$, I estimate the following specification:
+$\frac{1}{s}\ln\Big{(}\frac{y_{i,t+s}}{y_{i,t}}\Big{)}=\beta_{0}-\Big{(}\frac{1-\exp^{\beta s}}{s}\Big{)}\ln(y_{i,t})+\epsilon_{i,t+s}.$ (1)
+Here, $y_{i,t}$ refers to the income of country $i$ in year $t$, $y_{i,t+s}$ refers to the income of country $i$ in year $t+s$, and $\epsilon_{i,t+s}$ captures the shocks to the annualized growth rate between time periods $t$ and $t+s$.
 I assume that these shocks are independently and identically distributed (iid) with a normal distribution.
+I estimatate specification 1 using non-linear least square with heteroskedaticity-robust standard errors.
 I estimate separate regressions for the periods 1980-2000 and 2000-2019.
@@
 Additionally, note that the regression in equation 1 includes no country-level controls, meaning that the $\beta$ identified here reflects the rate of unconditional convergence.
-In both cases, I fix the base-year income at its 1980 level.
-Supplementary Appendix provides additional analysis, running the convergence regression with 1960 as the base year.
 Table 1: Beta Convergence for 1980-2000 and 2000-2019.
@@
 1980-2000 2000-2019 1980-2000 2000-2019
-$\beta$ 0.005** -0.008*** -0.001 -0.015***
+$\beta$ 0.0047** -0.0062*** 0.0006 -0.0150***
 (0.001) (0.002) (0.001) (0.003)
-N 131 131 91 91
-R2 0.046 0.099 0.001 0.238
+N 131 155 86 115
 Standard errors in parentheses
 * $p<0.10$, ** $p<0.05$, *** $p<0.01$
-Table 1 presents the estimates for the rate of $\beta-$convergence.
-Columns (1) and (2) report the estimates for the full balanced panel of countries, while Columns (3) and (4) show results for the set of countries excluding Sub-Saharan Africa.
-In Column (1), the convergence coefficient is positive and significant for the full sample of countries during the 1980-2000 period.
-This indicates that rich countries were initially growing faster than poor countries, leading to increased income divergence.
-However, for countries outside Sub-Saharan Africa, the convergence coefficient is statistically indistinguishable from zero during the same period, providing no evidence of convergence in either sample.
+Table 1 presents estimates of the rate of $\beta$-convergence.
+Columns (1) and (2) report results for the full balanced panel, whereas columns (3) and (4) restrict the sample by excluding Sub-Saharan African economies.
+Over the 1980-2000 period, the coefficient in column (1) is positive and statistically significant, indicating that wealthier countries grew faster than poorer ones and thereby increasing income divergence.
+The coefficient in column (3) does not differ significantly from zero, showing that even when Sub-Saharan Africa is excluded, there was no convergence over the 1980-2000 period.
 In contrast, the rate of convergence has seen a remarkable reversal in the past two decades.
-In the full sample, initially poor countries have been catching up to initially rich countries at an annual rate of 0.8%.
-For countries outside Sub-Saharan Africa, the catch-up rate has been even faster, at 1.5% annually.
-For comparison, [18] found a the rate of convergence of 0.425% for the full sample of countries and 1.175% outside Sub-Saharan Africa between 2000-2019 with 2000 as the base year.
-[15] don’t report the rate of convergence outside Sub-Saharan Africa but do find the rate of convergence to be 0.7% in broad sample of countries between 2005-2015 with 2005 as the base year.
+In the full balanced panel, initially poorer economies have closed their income gap at an annual rate of 0.62 percent.
+When Sub-Saharan African countries are excluded, the catch-up rate rises to 1.5 percent per year.
+By comparison, [17] report a 0.425 percent convergence rate for 2000-2019 in their full sample and 1.35 percent outside Sub-Saharan Africa, while [14] document a 0.7 percent rate for 2005–2015 (base year 2005) in a broad country panel.
+The somewhat higher coefficient estimated here reflects differences in sample composition and output measurement: [17] employ expenditure-side real GDP (rgdpe) and omit countries with populations below one million, whereas this analysis uses output-side real GDP (rgdpo) and excludes economies with fewer than 0.2 million inhabitants, in line with [14].
+Table 3 presents the convergence coefficients across alternative Penn World Table output series for both samples.
+The unconditional convergence coefficient for the post 2000 period outside SSA is also close to the rate of conditional convergence found in recent literature.
 [1] estimates the rate of conditional convergence to be 1.7% regressing 5-year growth rates on lagged income level and country level-time varying controls.
-The unconditional convergence rate of 1.5% observed here is substantial in comparison.
-This rate of $\beta-$convergence implies a half-life of 46 years, meaning that, starting from 2019, poor countries outside Sub-Saharan Africa would close half the income gap with developed countries by 2065, assuming all countries share the same steady-state income level.
-[^5] The half-life is calculated as $\tau=-\ln(2)/\ln(1-\beta)$.
-Moreover, nearly 24% of the variation in growth rates across countries outside Sub-Saharan Africa can be explained by their income levels in 1980.
-Macro growth regressions typically suffer from low explanatory power since there are a lot of variables determining future growth trajectories not included in the model.
-The relatively high $R^{2}$ achieved here without the inclusion of additional covariates, suggests that a country’s position in the income distribution is a strong proxy for a range of other factors affecting subsequent growth rates.
-Given that most of the convergence post-2000 is observed outside Sub-Saharan Africa, and that data for countries outside Sub-Saharan Africa is generally more reliable [21], the remainder of the paper focuses on this subset of countries.
-Results for the full set of countries are provided in the Supplementary Appendix.
-4 Trends in cross country income differences
-This section documents that the decline in overall dispersion of income among counties outside Sub-Saharan Africa started to decline post 2000.
-Table 2 presents the evolution of income dispersion across countries outside Sub-Saharan Africa over time, using various measures.
-Columns 1, 2, and 3 display the P90/P10, P90/P50, and P50/P10 percentile ratios of GDP per capita, which capture different aspects of income inequality across the distribution.
-Column 4 reports the variance of the log of GDP per capita, and Column 5 shows the income ratio between the top 5 and bottom 5 countries.
-The data reveal that all measures of income dispersion increased between 1980 and 1990.
-While the P90/P10 and P50/P10 ratios began to decline after 1990, the most notable trend is the consistent decrease in all measures of income inequality starting from 2000.
-By 2019, each measure of income dispersion reached its lowest point in the last 40 years.
+The unconditional convergence rate of 1.5% observed here for countries outside SSA is substantial in comparison.
+This rate of $\beta-$convergence rate of 1.5% implies a half-life of 53 years, meaning that, starting from 2020, poor countries outside Sub-Saharan Africa would close half the income gap with developed countries by 2073, assuming all countries share the same steady-state income level.
+[^4] The half-life is calculated as $\tau=-\ln(2)/\ln\Big{(}1-\frac{1-\exp^{\beta s}}{s}\Big{)}$.
 Table 2: Income Differences across years (outside SSA)
 Year P90/P10 P90/P50 P50/P10 Var(log(GDPpc)) Income Ratio
+Outside SSA
 1980 15.75 3.22 4.89 1.08 35.29
@@
 2019 9.97 2.73 3.64 0.81 24.78
-Given the significant decline in cross-country income dispersion, the next section conducts an accounting exercise to decompose this convergence into parts attributable to convergence in TFP, the capital-output ratio, and human capital.
+All Countries
+1980 18.56 5.30 3.50 1.15 18.49
+1990 24.95 5.55 4.50 1.37 23.26
+2000 31.95 6.88 4.64 1.62 29.53
+2010 27.63 4.51 6.12 1.54 30.71
+2019 25.06 4.41 5.69 1.47 27.17
+Beta-convergence is necessary but not sufficient for sigma-convergence [22, 3].
+[^5] Sigma convergence refers to a decline over time in the cross-sectional dispersion of per-capita income across economies.
+Consequently, even if initially poorer economies grow more rapidly than their richer counterparts, the cross-sectional dispersion of per-capita income need not decline.
+Table 2 reports five measures of income dispersion-namely the P90/P10, P90/P50 and P50/P10 percentile ratios of GDP per capita; the variance of log GDP per capita; and the ratio of the top five to bottom five incomes-for each decade from 1980 to 2019.
+The table enables a clear assessment of whether-and when-sigma-convergence accompanied the episodes of unconditional convergence identified earlier.
+For countries outside SSA, income dispersion widened during the 1980s—by 1990 the P90/P10 ratio had risen to nearly 20 and the variance of log GDP per capita to above 1.1.
+Since 2000, every measure has fallen steadily: by 2019 the P90/P10 ratio had roughly halved to below 10, and the variance of log GDP per capita declined to 0.8, reaching its lowest level in four decades.
+Hence, after 2000, initially poorer non-SSA economies not only grew faster than their richer counterparts but also saw a marked reduction in overall income dispersion—demonstrating both beta- and sigma-convergence in that subsample.
+For the full country sample, dispersion rose sharply through the 1980s and 1990s, with the P90/P10 ratio climbing from 18.6 in 1980 to nearly 32.0 in 2000 before easing to 25.1 by 2019.
+Likewise, the variance of log GDP per capita peaked at 1.62 in 2000 and then fell modestly to 1.47 by 2019.
+By contrast, the P50/P10 ratio rose steadily from 3.50 in 1980 to 5.69 in 2019, signaling a widening gap between middle-income and poorest countries.
+These trends imply only limited sigma-convergence: although top-to-bottom dispersion declined somewhat after 2000, overall income inequality in 2019 remains well above its 1980 level.
+Table 2 demonstrates that, although economies outside Sub-Saharan African achieved both beta- and sigma-convergence after 2000, many Sub-Saharan African countries have continued to stagnate.
+Since most of the observed convergence-and the most reliable data-resides outside SSA [21], the subsequent analysis focuses on non-SSA economies.
+Full-sample results appear in the Appendix.
 4 Accounting for Decline in Income Dispersion
@@
 Unlike many standard approaches in the literature, which assume a uniform capital share parameter of one-third, this accounting exercise based on percentile ratios allows the capital share of income to vary continuously with the income percentile, as observed in PWT10.01.
-The accounting exercise using variance as a measure of income dispersion is presented in the supplementary appendix.
-Let $p$ denote the income percentile.
-Suppose that per capita output ($y$) is a function of physical capital ($k$), human capital ($h$), and a residual term ($A$).
-[^6] It is common to interpret the residual term ($A$) as total factor productivity (TFP), but more broadly, it includes both observable and unobservable factors that affect output beyond the inputs.
-The output production function, parameterized by $p$, takes the following form:
-$y(p)$ $=A(p)(k(p))^{\alpha(p)}(h(p))^{1-\alpha(p)}$
-This is the standard Cobb-Douglas production function, with the modification that the capital share ($\alpha$) is allowed to vary across the income distribution.
-This flexibility is motivated by the large variation in capital share across countries and over time.
-Figure 1 shows the capital share along the income distribution, as computed from PWT10.01.
-The capital share is not only heterogeneous but higher than what is typically assumed in the literature.
-The production function can be rearranged and written in intensive form as follows:
-$y(p)=A(p)^{\frac{1}{1-\alpha(p)}}(\frac{k}{y}(p))^{\frac{\alpha(p)}{1-\alpha(p)}}(h(p)).$
-Writing the production function in this form reduces the effect of differences in per capita capital ($k$), since per capita capital is allowed to adjust endogenously in response to changes in total factor productivity ($A$).
-[^7] As argued in [14], under firm optimization, equalization of the marginal product of capital with the rental rate means that an increase in total factor productivity (TFP) leads to a mechanical increase in capital, assuming a constant rental rate.
-Since my goal is to separately identify the contributions of total factor productivity (TFP) growth and capital accumulation to output growth, I prefer to use the production function written in this form.
-Given this functional form, the income difference between percentiles $p_{0}$ and $p_{n}$ can be decomposed as follows:
+The accounting exercise using variance as a measure of income dispersion is presented in the appendix.
+Let $p$ denote the income percentile. Suppose per-worker output is generated by a constant-returns-to-scale technology
+$y(p)=F\bigl{(}A(p),k(p),h(p)\bigr{)},$
+where $A(p)$ is Hicks-neutral productivity.
+[^6] It is common to interpret $A(p)$ as total factor productivity (TFP), but more broadly it includes both observable and unobservable factors that affect output beyond the inputs.
+As argued in [13], under firm optimization, equalization of the marginal product of capital with the rental rate implies that TFP growth mechanically raises the capital stock under a constant rental rate; to separate TFP and capital-accumulation effects in output growth, the production function is written in intensive form.
+Then the difference in log-output between percentiles $p_{0}$ and $p_{n}$ admits the decomposition
 […]
-Here, the second and third terms capture the role of inputs, while the residual term ensures that the accounting identity holds.
-Appendix A provides a detailed explanation of how each component is computed.
-I report the results from the accounting exercise for three percentile ratios-P90/P10, P90/P50, and P50/P10-covering countries outside Sub-Saharan Africa for the years 2000-2019, and 1980-2019.
-These ratios capture the overall trend in convergence outside Sub-Saharan Africa across the income distribution.
-Table 3: Accounting for Income Convergence outside SSA
-Percentile Ratio $\Delta_{t}\log(\frac{y(p_{n})}{y(p_{0})})$ Cont. Cont. Cont.
-(%) A (%) $\frac{K}{Y}$ (%) $h$ (%)
-A: 1980-2000
-P90/P10 8 487 -225 -175
-P90/P50 18 194 -67 -22
-P50/P10 -10 -40 60 90
-B: 2000-2019
-P90/P10 -54 44 43 11
-P90/P50 -34 35 44 26
-P50/P10 -20 65 40 -15
-C: 1980-2019
-P90/P10 -46 -33 89 43
-P90/P50 -16 -144 169 81
-P50/P10 -30 30 47 20
-Table 3 presents the results of the accounting exercise, showing changes in income percentile ratios over three periods: 1980-2000, 2000-2019, and 1980-2019.
-Positive values in the changes of income percentile ratios signify income divergence, indicating that income inequality between countries has increased, while negative values denote income convergence, meaning that income inequality has decreased.
-The table breaks down the contributions to these changes from three factors: total factor productivity (TFP) differences, capital-output ratios, and human capital differences.
-Each row of the table details these contributions across three key percentile ratios-P90/P10, P90/P50, and P50/P10-which represent disparities between the rich (90th percentile) and the poor (10th percentile), the rich and the middle-income group (50th percentile), and the middle-income group and the poor, respectively.
-The sign of each percentage contribution indicates the direction of its contribution relative to the overall change: a positive percentage means the factor’s change aligns with the total change, while a negative percentage means it opposes the total change.
-For example, during a period of income divergence (a positive change in the percentile ratio), a positive contribution from TFP differences implies that TFP is also diverged during the same period.
-Similarly, if the percentile ratio is converging (a negative change), a positive contribution from TFP differences indicates that TFP is also converged during the same period.
-Panel A of Table 3 shows that between 1980 and 2000, the P90/P10 and P90/P50 income ratios diverged, while the P50/P10 ratio converged.
-Although the capital-output ratio and human capital converged between rich and poor economies and between rich and middle-income economies, TFP diverged among these groups, leading to overall income divergence.
-Notably, the income divergence would have been even greater if inputs had not converged.
-In contrast, the income difference between middle-income and poor economies converged during this period.
-The convergence of inputs accounts entirely for the income convergence between middle-income and poor economies, as TFP diverged between these groups during this time.
-Panel B reveals that, unlike the 1980-2000 period, income differences converged across the income distribution during 2000-2019.
-This shift was due to TFP diverging in the earlier period to converging in the later period.
-More importantly, the reduction in the gaps in the capital-output ratio significantly contributed to the narrowing of income differences during 2000-2019.
-Specifically, the P90/P10 income ratio declined by 54%, with the decline in the capital-output ratio accounting for 43% of this reduction.
-Similarly, the P90/P50 ratio declined by 34%, and the P50/P10 ratio declined by 20%.
-The decline in the capital-output ratio explains 44% of the reduction in P90/P50 differences and 40% of the reduction in P50/P10 differences.
-These findings highlight the critical role of physical capital in driving convergence during this period.
-When changes in the capital-output ratio are combined with improvements in human capital, the total contribution of inputs becomes even more pronounced: inputs account for 52% of the reduction in P90/P10 disparities, 65% of the reduction in P90/P50 disparities, and 35% of the reduction in P50/P10 disparities.
-The overall decline in input dispersion accounts for more than half of the reduction in output disparities between the richest (P90) and poorest (P10) economies, as well as between the richest and middle-income (P50) economies.
-Panel C presents the findings for the extended period from 1980 to 2019.
-During this timeframe, convergence in the capital-output ratio was the primary driver of narrowing income disparities.
-The P90/P10, P90/P50, and P50/P10 income ratios declined by 46%, 16%, and 30%, respectively.
-The reduction in the capital-output ratio accounts for 89% of the decrease in P90/P10 income differences, while TFP differences contributed negatively with -33%, indicating divergence.
-For the P90/P50 ratio, the capital-output ratio explains a striking 169% of the reduction, offsetting the negative contribution from the TFP divergence of -144%.
-For the P50/P10 ratio, the capital-output ratio accounts for 47% of the convergence, with TFP contributing 30%.
-When changes in the capital-output ratio are combined with improvements in human capital, inputs account for 132% of the reduction in P90/P10 disparities, 250% of the reduction in P90/P50 disparities, and 77% of the reduction in P50/P10 disparities.
-These results imply that, had it not been for the significant convergence in the inputs, especially the capital-output ratio, the overall income convergence between the richest and poorer countries would have been substantially lower or potentially nonexistent due to TFP divergence during this period.
-5.1 Accounting with constant capital income share
-Table 4: Accounting for Income Convergence outside SSA ($\alpha=1/3$)
-Percentile Ratio $\Delta_{t}\log(\frac{y(p_{n})}{y(p_{0})})$ Cont. Cont. Cont.
-(%) A (%) $\frac{K}{Y}$ (%) $h$ (%)
-A: 1980-2000
-P90/P10 8 412 -137 -175
-P90/P50 18 178 -44 -22
-P50/P10 -10 -10 30 90
-B: 2000-2019
-P90/P10 -54 65 24 11
-P90/P50 -34 53 24 26
-P50/P10 -20 85 25 -15
-C: 1980-2019
-P90/P10 -46 4 52 43
-P90/P50 -16 -88 100 81
-P50/P10 -30 53 27 20
-Table 4 presents the main accounting result assuming a constant capital share of one-third.
-Panel A of Table 4 shows that between 1980 and 2000, the divergence of TFP among rich and poor economies, as well as between poor and middle income economies still accounts for all the divergence in income ratios.
-This suggests that for studies focused on this time period, such as [14], the assumption of a constant capital share does not make much of a difference.
-Panel B of Table 4, shows that between 2000 and 2019, TFP accounts for the majority of convergence among rich and poor, middle and poor, and rich and middle income economies.
-This contrasts with the results in Panel B of Table 3, where inputs accounts for majority of the convergence.
-Therefore, for the period post-2000, the choice of capital income share becomes crucial.
-When heterogeneity of capital income shares is incorporated into the growth accounting exercise, capital alone accounts for 43%, 44%, and 40% of the convergence in P90/P10, P90/P50, and P50/P10 ratios between 2000 and 2019.
-In comparison, with the assumption of a constant capital income share of one-third, the convergence of capital accounts for only 24%, 24%, and 25% of the convergence, respectively.
-Hence, not incorporating heterogeneity in capital income shares during this time period would lead to the incorrect conclusion that TFP remains the dominant factor driving growth and convergence.
-Panel C of Table 4 shows that for the extended period from 1980 to 2019, with a constant one-third capital income share, convergence of capital accounts for 52%, and 27% of the convergence in P90/P10 and P50/P10 ratios, respectively, compared to 89% and 47% when heterogeneity of capital income shares is included.
-Capital still accounts for all the convergence in the P90/P50 ratio in both cases.
-Therefore, even in the long run, the assumption of a constant capital income share significantly reduces the role of capital in driving income convergence between rich and poor, and between middle-income and poor economies.
-Overall, the findings indicate that the convergence of inputs-particularly physical capital-accounts for all the income convergence between rich and poor economies during 1980-2019, and more than half of the convergence during 2000-2019 outside Sub-Saharan Africa.
-Moreover, the choice of capital income share is important.
-Assuming a constant capital share of one-third reduces the role of capital in driving convergence by approximately half.
+where $\alpha(p)$ denotes the capital share of income at percentile $p$.
+Here, the first term captures the contribution of TFP, the second term captures the contribution of capital-output ratio, and the third term captures the contribution of human capital.
+Appendix A provides the full derivations of each elasticity under CRS.
+I report the results from the decomposition exercise for the percentile ratio-P90/P10- covering countries outside Sub-Saharan Africa for the time periods: 1980-2000, 2000-2019.
+The decomposition results for P90/P50, P50/P10, as well as for full sample of countries are reported in Appendix.
+[^7] I refer P90, P50, P10 as rich, middle, poor economies respectively.
+Decomposition Results: Convergence outside SSA
+Figure 1: Decomposition Exercise (excluding SSA): P90/P10
+Figure 1 illustrates the decomposition of the P90/P10 income ratio over two distinct time periods: 1980-2000 (blue bars), and 2000-2019 (red bars).
+The vertical axis shows the change in values for each period, while the horizontal axis shows the overall change in the log P90/P10 ratio followed by the contributions of TFP, the capital-output ratio, and human capital to that change, as derived in equation (2).
+Convergence in both physical and human capital inputs was the principal force shaping the evolution of the P90/P10 income ratio.
+Between 1980 and 2000, the P90/P10 income ratio increased by 8 percentage points, reflecting a period of divergence between rich and poor countries.
+Before 2000, TFP divergence alone would have driven the ratio up by roughly 0.39, but simultaneous convergence in physical capital and human capital reduced the ratio by about 0.17 and 0.14, respectively.
+From 2000 to 2019 the pattern reversed.
+The log P90/P10 ratio fell by $0.54$.
+Convergence in the capital-output ratio contributed $0.24$ of this reduction (about 44%), while human-capital convergence contributed $0.06$ (10%).
+During this interval TFP shifted from divergence to convergence, contributing $0.25$ (46%) to overall convergence.
+The shift in total factor productivity (TFP) dynamics—from divergence before 2000 to convergence thereafter—played a pivotal role in driving post-2000 income convergence.
+In the absence of TFP convergence, the extent of overall convergence during this period would have been minimal.
+Nevertheless, in quantitative terms, the contribution of TFP to post-2000 convergence remains smaller than that of input accumulation, accounting for 46% compared to 54%.
+This contrast is even more pronounced when examining the earlier period from 1980 to 2000.
+Note that the overall change in the P90/P10 ratio from 1980 to 2019 is the sum of the changes over 1980-2000 and 2000-2019, as expressed by the additive property of logarithms.
+Over the full horizon 1980-2019 the log P90/P10 ratio declined by $0.46$.
+Cumulatively, the capital-output ratio accounted for $0.41$ of the decline (89%) and human capital for $0.19$ (41%), whereas persistent TFP divergence added back roughly $0.14$.
+Despite the continued divergence in TFP, there was substantial convergence in income over the period 1980-2019.
+This implies that convergence in physical and human capital inputs fully accounts for the observed income convergence during this period.
+The Appendix reports analogous decompositions for the P90/P50 and P50/P10 ratios, respectively.
+For the P90/P50 ratio input convergence accounts for 66% of the decline observed in 2000-2019 and explains the entire convergence over the full 1980-2019 horizon.
+In the case of the P50/P10 ratio inputs contribute less—32% of the convergence during 2000-2019 and 70% over 1980-2019.
+In both decompositions the behavior of total factor productivity mirrors that documented for the P90/P10 ratio, diverging in 1980-2000 and converging in 2000-2019.
+Taken together, the evidence indicates that poorer economies outside Sub-Saharan Africa are catching up not only with rich but also with middle-income countries.
+This long-run convergence—particularly after 2000, and more importantly over the full period from 1980 to 2019—is driven primarily by the convergence of physical and human capital inputs.
+Decomposition Results: Stagnation of SSA
+Figure 2: Decomposition Exercise (full sample): P90/P10
+Figure 2 presents the decomposition results for the full sample of countries, including those in Sub-Saharan Africa (SSA).
+It is important to note that, as of 2000, economies below the P10 income level were predominantly from Sub-Saharan Africa.
+Therefore, this decomposition provides insights into the stagnation of poor economies in SSA relative to those outside the region.
+Similar to poor countries outside SSA, economies in SSA experienced divergence in income during 1980-2000 and convergence toward richer countries after 2000.
+However, the degree of divergence before 2000 was substantially larger within SSA (0.54 compared to 0.08), and the extent of convergence since 2000 was correspondingly much smaller (0.27 compared to 0.54).
+The larger divergence before 2000 for SSA economies is predominantly driven by more pronounced divergence in total factor productivity (TFP): 0.71 compared to 0.39 for non-SSA countries.
+Lower convergence in human capital also contributed: 0.03 for SSA compared to 0.14 outside SSA—but its role was relatively modest in comparison to TFP.
+Convergence in physical capital, by contrast, was broadly similar across the two groups, at 0.13 for SSA and 0.17 for non-SSA economies.
+In the post-2000 period, convergence in TFP (0.22 compared to 0.25) and in human capital (0.05 compared to 0.06) were of similar magnitudes between SSA and non-SSA poor economies.
+The key distinction lies in physical capital: the absence of convergence in physical capital among SSA economies entirely accounts for the lower extent of output convergence relative to poor countries outside the region.
+Unlike poor economies outside SSA, SSA did not exhibit any meaningful convergence in the capital-output ratio post-2000, and only negligible convergence in human capital.
+While TFP did converge post-2000, the lack of convergence in inputs meant that the forces of convergence were insufficient to offset the TFP divergence that occurred prior to 2000.
+As a result, between 1980 and 2019, SSA economies fell further behind the rich economies due to TFP divergence in the early period and the absence of input convergence across the full time span.
+Further evidence of the lack of input convergence, particularly in capital, is presented in Figure 3, which illustrates the mean capital-output ratio, weighted by population size, across different regions.
+While the capital-output ratio has continued to rise in all regions outside Sub-Saharan Africa, SSA is the only region to exhibit a decline in its capital-output ratio post-2000.
+This decline likely reflects the region’s struggles with capital accumulation—challenges that may be linked to political instability, economic uncertainty, and capital flight, as discussed in [16].
+Figure 3: Capital-Output Ratio across regions
+Decomposition with constant capital income share
+Figure 4: Variation of alpha over income percentile.
+Figure 4 illustrates the capital share of income across countries at different percentiles of the income distribution for various years.
+For the full sample of countries included in this analysis, PWT shows an average capital share of income of 0.46, which significantly exceeds the commonly assumed value of 1/3 in the literature.
+Moreover, the capital share of income follows an inverse U-shaped pattern across the income distribution, with middle-income countries having the highest capital income share, followed by poor countries, and high-income countries having the lowest capital income share.
+Only the richest countries have capital income shares close to 1/3, suggesting that using a constant capital income share may not be appropriate for countries at other points in the income distribution.
+The decomposition approach in the current paper accommodates this heterogeneity, allowing the capital income share to vary along the income distribution.
+Additionally, over time, countries across the income distribution have experienced an increase in the capital share of income, which indicates a corresponding decline in labor share.
+The recent work by [12] examines this global decline in labor share and attributes much of the trend to cheaper information and communication technology (ICT) capital.
+Because capital-income shares differ systematically across the income distribution, any growth-accounting exercise must incorporate this heterogeneity.
+Figure 5 reports-for comparison—the results obtained when a constant capital share of one-third is imposed.
+Figure 5: Decomposition Exercise (excluding SSA, alpha=1/3): P90/P10
+During the 1980-2000 period cross-country income divergence is still driven overwhelmingly by the widening gap in total factor productivity, with the convergence of physical and human capital acting in the opposite direction and therefore dampening the TFP effect.
+Because TFP remains the dominant force in the accounting, studies that concentrate on the pre-2000 era—such as [13]—are not materially affected by whether one allows the capital share of income to vary across countries or, instead, fixes it at the conventional one-third benchmark.
+However, for the post-2000 period, the choice of capital income share becomes important.
+With a constant one-third share, the capital-output ratio accounts for only 23% of the decline in the P90/P10 ratio, compared to 44% when heterogeneity in capital income shares is incorporated.
+This difference implies that ignoring heterogeneity may lead to the erroneous conclusion that TFP remains the dominant factor in driving growth and convergence.
+Over the full period from 1980 to 2019, the constant one-third assumption results in capital convergence explaining 52% of the overall convergence, as opposed to 89% when capital income shares vary by income.
+Thus, even in the long run, assuming a constant capital share significantly underestimates the role of capital in driving income convergence between rich and poor economies.
+The Appendix further presents the decomposition results for the P90/P50 and P50/P10 ratios under the constant one-third assumption, showing a quantitatively reduced role of capital in driving convergence compared to the case when capital income shares are allowed to vary.
+The results remains unchanged when looking at full sample of countries.
 5 Measurement of Inputs and Capital Share of Income
@@
 If returns to schooling, the quality of education, or returns to experience are also converging across economies, then the contribution of human capital presented here is a lower bound.
-Measurement of Capital Share of Income
+Capital Share of Income
 The capital share of income ($\alpha$) in PWT10.01 is derived using the following relationship:
@@
 Whenever mixed income data is available, PWT10.01 applies the second adjustment.
-In the absence of mixed-income data, I use the minimum of the third and fourth adjustments.
-Figure 1: Variation of alpha over income percentile.
-Figure 1 illustrates the capital share of income across countries at different percentiles of the income distribution for various years.
-For the sub-sample of countries included in this analysis, PWT shows an average capital share of income of 0.46, which significantly exceeds the commonly assumed value of 1/3 in the literature.
-Moreover, the capital share of income follows an inverse U-shaped pattern across the income distribution, with middle-income countries having the highest capital income share, followed by poor countries, and high-income countries having the lowest capital income share.
-Only the richest countries have capital income shares close to 1/3, suggesting that using a constant capital income share may not be appropriate for countries at other points in the income distribution.
-The decomposition approach in the current paper accommodates this heterogeneity, allowing the capital income share to vary along the income distribution.
-Additionally,, over time, countries across the income distribution have experienced an increase in the capital share of income, which indicates a corresponding decline in labor share.
-The recent work by [13] examines this global decline in labor share and attributes much of the trend to cheaper information and communication technology (ICT) capital.
-The growing importance of the capital share of income reinforces the hypothesis that the convergence of capital has played a substantial role in the convergence of output post-2000.
+In the absence of mixed-income data, PWT10.01 uses the minimum of the third and fourth adjustments.
 6 Conclusion
-This study revisited the dynamics of cross-country income convergence using the most recent data from the Penn World Tables (PWT 10.01).
-Roughly 70% of cross-country empirical work relies on the Penn World Tables [11].
-The updated PWT data no longer support some of the basic growth accounting facts that underpinned earlier studies.
+This study revisited the dynamics of cross-country income convergence using the most recent data from the Penn World Tables (PWT 10.01).999 Roughly 70% of cross-country empirical work relies on the Penn World Tables [10].
 Contrary to the traditional view that emphasizes non-convergence and the limited role of factor inputs, the analysis reveals a significant shift in global income patterns since 2000.
@@
 Since input accumulation, rather than TFP growth, primarily drives convergence, policies aimed at removing barriers to the accumulation of physical and human capital in poorer economies may yield more substantial growth dividends than those focused solely on TFP enhancements.
+[…]
+Acknowledgements
+I am grateful to Anmol Bhandari, Alessandra Fogli, Hannes Malmberg, Todd Schoellman, and Kjetil Storesletten for their advice and support.
+I would also like to thank Honey Batra, Francisco Bullano, G V A Dharanan, Jason Hall, Sang Min Lee, Ruyan Liao, Martin Garcia Vazquez, and Teerat Wongrattanapiboon for useful comments.
 Appendix A Mathematical Derivations
 A.1 Generalized Decomposition
-Let $p$ denote percentile and output be some arbitrary function if inputs, TFP parametrized by $p$.
-$y(p)=F(A(p),k/y(p),h(p)).$
-Then,
+Let $p$ denote the percentile index, and consider the per-worker output function
+$y(p)=F\bigl{(}A(p),\,k(p),\,h(p)\bigr{)}=F\bigl{(}A(p),\,(k(p)/y(p))\,y(p),\,h(p)\bigr{)},$ (3)
+where $\kappa(p)=k(p)/y(p)$. Total differentiation of (3) with respect to $p$ gives
+$\frac{dy}{dp}$ $=F_{A}\bigl{(}A,\kappa y,h\bigr{)}\,\frac{dA}{dp}+F_{k}\bigl{(}A,\kappa y,h\bigr{)}\,\frac{d(\kappa y)}{dp}+F_{h}\bigl{(}A,\kappa y,h\bigr{)}\,\frac{dh}{dp},$ (4)
+$\frac{d(\kappa y)}{dp}$ $=\kappa\,\frac{dy}{dp}+y\,\frac{d\kappa}{dp}.$ (5)
+Rearranging yields
+$\bigl{[}1-F_{k}(A,\kappa y,h)\,\kappa\bigr{]}\,\frac{dy}{dp}=F_{A}(A,\kappa y,h)\,\frac{dA}{dp}+F_{k}(A,\kappa y,h)\,y\,\frac{d\kappa}{dp}+F_{h}(A,\kappa y,h)\,\frac{dh}{dp}.$ (6)
+Dividing by $y$ gives the differential form
+$\frac{d\ln y}{dp}=\underbrace{\frac{F_{A}}{y\,\bigl{(}1-F_{k}\,\kappa\bigr{)}}\,\frac{dA}{dp}}_{\mathcal{E}_{A}(p)}+\underbrace{\frac{F_{k}\,\kappa}{1-F_{k}\,\kappa}\,\frac{d\ln\kappa}{dp}}_{\mathcal{E}_{k/y}(p)}+\underbrace{\frac{F_{h}\,h}{y\,\bigl{(}1-F_{k}\,\kappa\bigr{)}}\,\frac{d\ln h}{dp}}_{\mathcal{E}_{h}(p)}.$ (7)
+Integration from $p_{0}$ to $p_{n}$ yields the generalized decomposition
 […]
-$\log(y(p_{n}))-\log(y(p_{0}))$ $=\int_{p_{0}}^{p_{n}}[\frac{d\log F(.)}{d\log A(p)}\frac{d\log A(p)}{dp}+\frac{d\log F(.)}{d\log k/y(p)}\frac{d\log k/y(p)}{dp}+\frac{d\log F(.)}{d\log h(p)}\frac{d\log h(p)}{dp}]$
+For a CRS technology in $(k,h)$, Euler’s theorem implies
+$F_{k}(A,k,h)\,k+F_{h}(A,k,h)\,h=y,\quad\alpha(p)=\frac{F_{k}(A,k,h)\,k}{y}.$ (8)
+Hence
+$F_{k}(A,k,h)\,\kappa=\alpha(p),\quad\frac{F_{h}(A,k,h)\,h}{y}=1-\alpha(p).$
+Using these identities, the three elasticities in (7) become
+$\mathcal{E}_{A}(p)$ $=\frac{F_{A}}{y\,(1-\alpha(p))}\,\frac{dA}{dp}=\frac{1}{1-\alpha(p)}\,\frac{d\ln A(p)}{dp},$
+$\mathcal{E}_{k/y}(p)$ $=\frac{\alpha(p)}{1-\alpha(p)}\,\frac{d\ln(k/y)(p)}{dp},$
+$\mathcal{E}_{h}(p)$ $=\frac{1-\alpha(p)}{1-\alpha(p)}\,\frac{d\ln h(p)}{dp}=\frac{d\ln h(p)}{dp}.$
+Substituting back into the integrated form yields the final result
 […]
-where
-$\mathcal{E}_{A}(p)$ $=\frac{d\log F}{d\log A}\frac{d\log A}{dp}$
-$\mathcal{E}_{k/y}(p)$ $=\frac{d\log F}{d\log(k/y)}\frac{d\log(k/y)}{dp}$
-$\mathcal{E}_{h}(p)$ $=\frac{d\log F}{d\log(h)}\frac{d\log(h)}{dp}$
-Now suppose that the output production has the following functional form:
-$y(p)$ $=A(p)(k(p))^{\alpha(p)}(h(p))^{1-\alpha(p)}$
-$=A(p)^{\frac{1}{1-\alpha(p)}}(k/y(p))^{\frac{\alpha(p)}{1-\alpha(p)}}h(p)$
-In this case:
-$\mathcal{E}_{A}(p)$ $=\frac{1}{1-\alpha(p)}\frac{d\log A}{dp}$
-$\mathcal{E}_{k/y}(p)$ $=\frac{\alpha(p)}{1-\alpha(p)}\frac{d\log(k/y)}{dp}$
-$\mathcal{E}_{h}(p)$ $=\frac{d\log(h)}{dp}$
 The contribution of $A$ can be calculated as a residual. The contribution of each of the inputs can then be evaluated as follows:
@@
 […]
-Note that if the elasticity parameter is constant along income distribution ( $\alpha(p)=\alpha$), the the generalized decomposition reduces to the usual Hall-Jones decomposition.
+Note that if the elasticity parameter is constant along income distribution ( $\alpha(p)=\alpha$), the generalized decomposition reduces to the usual Hall-Jones decomposition.
 $\ln(y(p_{n}))-\ln(y(p_{0}))$ $=\frac{1}{1-\alpha}\ln(\frac{A(p_{n})}{A(p_{0})})+\frac{\alpha}{1-\alpha}\ln(\frac{k/y(p_{n})}{k/y(p_{0})})+\ln(\frac{h(p_{n})}{h(p_{0})}).$
 Appendix B Robustness and Additional Exercises
-B.1 Unconditional Convergence with 1960 Base Year
-Table 5: Beta Convergence for 1980-2000 and 2000-2019.
-All Countries Outside SSA
-(1) (2) (3) (4)
-1980-2000 2000-2019 1980-2000 2000-2019
-$\beta$ 0.002 --0.009∗∗∗ --0.004 --0.014∗∗∗
-(0.003) (0.003) (0.002) (0.004)
-N 107 107 72 72
-R2 0.007 0.090 0.034 0.157
-Standard errors in parentheses
-* $p<0.10$, ** $p<0.05$, *** $p<0.01$
-Table 5 presents the results of $\beta$-convergence regressions, using 1960 as the base year for initial income levels.
-The key objective here is to determine whether the choice of base year affects the conclusions regarding income convergence.
-The results confirm that the findings are robust to the use of a different base year.
-For both the full sample of countries and the sub-sample excluding Sub-Saharan Africa (SSA), there is no evidence of convergence between 1980 and 2000.
-In this period, the $\beta$ coefficient is close to zero and statistically insignificant in both samples, indicating no relationship between initial income levels and subsequent growth rates.
-In contrast, the period 2000-2019 shows strong evidence of convergence, with statistically significant negative $\beta$ coefficients for both the full sample and the sub-sample excluding SSA.
-The magnitude of convergence is similar to that found when using 1980 as the base year.
-For the full sample of countries, the $\beta$ coefficient is --0.009, while for countries outside SSA, it is --0.014.
-These results suggest that, post-2000, poorer countries have been catching up with richer countries at similar rates, regardless of whether 1960 or 1980 is used as the starting point for income levels.
-B.2 Robustness to different measures of GDP
-Table 6: Beta Convergence for 2000-2019 (Excluding SSA)
-RGDPO RGDPE CGDPO CGDPE
-$\beta$-0.015∗∗∗ --0.014∗∗∗ --0.015∗∗∗ --0.014∗∗∗
-(0.003) (0.003) (0.003) (0.003)
-N 91 91 91 91
-R2 0.238 0.221 0.250 0.221
-Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
-The Penn World Table (PWT) provides multiple measures of GDP, converted into common currency units, using different approaches to national income accounting.
-Two measures of real GDP are reported: one based on the output side (RGDPO) and another based on the expenditure side (RGDPE).
-Similarly, two measures of current-price GDP are available: the output-side measure (CGDPO) and the expenditure-side measure (CGDPE).
-Table 6 presents a robustness check of the $\beta$-convergence results for the period 2000-2019, using different measures of per capita GDP for countries outside Sub-Saharan Africa.
-Each column corresponds to a different measure of GDP used in the calculation of per capita income.
-The results indicate that the rate of convergence remains remarkably stable across all four GDP measures.
-In each case, the $\beta$ coefficient is negative and statistically significant at the 1% level, with convergence rates ranging between --0.014 and --0.015.
-This consistency suggests that the pattern of income convergence for countries outside Sub-Saharan Africa is robust to the choice of GDP measure, whether based on real output, real expenditure, or current-price accounting.
-The $R^{2}$ values also indicate a similar explanatory power across all specifications, with values ranging from 0.221 to 0.250.
-B.3 Income difference for all countries
-Table 7: Income Differences across Years (Full Sample)
-Year P90/P10 P90/P50 P50/P10 Var(log(GDPpc)) Income Ratio
-1980 18.56 5.30 3.50 1.15 18.49
-1990 24.95 5.55 4.50 1.37 23.26
-2000 31.95 6.88 4.64 1.62 29.53
-2010 27.63 4.51 6.12 1.54 30.71
-2019 25.06 4.41 5.69 1.47 27.17
+B.1 Robustness to different measures of GDP
+Table 3: Beta–Convergence Coefficients Across GDP Measures
+GDP Measure
+RGDPO RGDPE
+Panel A: All Countries
+Current sample $\beta$ –0.0062 –0.0052
+(SE) (0.0017) (0.0016)
+$N$ 155 155
+PSS sample $\beta$ –0.0052 –0.0043
+(SE) (0.0016) (0.0016)
+$N$ 124 124
+Panel B: Outside SSA
+Current sample $\beta$ –0.0150 –0.0128
+(SE) (0.0023) (0.0022)
+$N$ 115 115
+PSS sample $\beta$ –0.0155 –0.0135
+(SE) (0.0021) (0.0021)
+$N$ 90 90
+Notes: $\beta$ coefficients are estimated by nonlinear least squares as in (1), with robust standard errors in parentheses.
+RGDPO and RGDPE denote output-side and expenditure-side real GDP per capita at chained PPPs (million 2017 US\$) from PWT 10.01.
+PSS sample refers to the sample of country used in [17].
 […]
-Table 7 presents income differences across various measures for the full sample of countries, complementing the analysis of countries outside Sub-Saharan Africa (SSA) provided in the main text.
-The table reports several key indicators of income dispersion, including the P90/P10, P90/P50, and P50/P10 percentile ratios of GDP per capita, the variance of the log of GDP per capita, and the income ratio between the top 5 and bottom 5 of countries.
-A key observation from this table is that the decline in income disparities post-2000 has been relatively modest for the full sample of countries, particularly compared to the more substantial reductions observed in countries outside SSA.
-For instance, the P90/P10 ratio fell from 31.95 in 2000 to 25.06 in 2019, indicating some reduction in inequality between the richest and poorest economies.
-However, this decline is less pronounced compared to the convergence seen outside SSA, as shown in the main text.
-Notably, the P50/P10 ratio increased from 4.64 in 2000 to 5.69 in 2019, reflecting a widening income gap between middle-income and poor economies.
-This increase is largely driven by the stagnation of many Sub-Saharan economies, which have struggled to keep pace with global growth.
-Consequently, while income disparities have narrowed somewhat at the top end of the distribution (as seen in the decline in the P90/P10 and P90/P50 ratios), the lower half of the income distribution has experienced greater divergence, especially between middle-income and poorer countries.
-B.4 Accounting for Income Convergence (Full Sample)
-Table 8: Accounting for Income Convergence (Full Sample)
-Year $\Delta_{t}\log(y(p_{n})/y(p_{0}))$ Contribution Contribution Contribution
-(%) A (%) K/Y (%) $h$ (%)
-Panel A: 1980-2000
-P90/P10 54 131 -24 -6
-P90/P50 15 247 -67 -87
-P50/P10 39 85 -10 23
-Panel B: 2000-2019
-P90/P10 -27 81 0 19
-P90/P50 -29 48 24 24
-P50/P10 3 -233 267 67
-Panel C: 1980-2019
-P90/P10 27 181 -48 -30
-P90/P50 -14 -164 121 143
-P50/P10 42 62 10 26
-Table 8 presents the accounting decomposition for income differences across the global income distribution for three periods: 1980-2000, 2000-2019, and 1980-2019.
-The analysis includes all economies, including Sub-Saharan Africa.
-Panel A of Table 8 shows that between 1980 and 2000, income disparities widened across the income distribution.
-The P90/P10 income ratio increased by 54%, indicating divergence between the richest and poorest countries.
-TFP differences contributed significantly to this divergence, accounting for 131% of the increase.
-Convergence in the capital-output ratio and human capital partially offset this effect, contributing -24% and -6%, respectively.
-Similarly, the P90/P50 ratio increased by 15%, with TFP differences contributing 247% to divergence, while capital-output ratios and human capital contributed negatively, offsetting some of the divergence.
-The P50/P10 ratio increased by 39%, mainly due to TFP differences (85%), with minimal offset from inputs.
-Panel B of Table 8 reveals that between 2000 and 2019, income disparities decreased between the richest and poorest countries.
-The P90/P10 income ratio declined by 27%, indicating convergence.
-TFP differences contributed 81% to this convergence, while human capital contributed 19%.
-The capital-output ratio had a negligible effect in this case.
-The P90/P50 ratio decreased by 29%, with TFP and inputs contributing equally to convergence.
-Notably, the P50/P10 ratio showed a slight divergence of 3%.
-This divergence was due to a negative contribution from TFP (-233%), but substantial convergence in the capital-output ratio (267%) and human capital (67%) more than compensated for this, indicating significant input-driven convergence among middle-income and poor countries.
-Panel C of Table 8 presents the findings for the entire period from 1980 to 2019.
-Over this timeframe, the P90/P10 income ratio increased by 27%, indicating overall divergence between the richest and poorest countries.
-TFP differences were the main driver, contributing 181% to this divergence.
-Convergence in the capital-output ratio (-48%) and human capital (-30%) partially mitigated the divergence caused by TFP.
-The P90/P50 ratio decreased by 14%, reflecting convergence between the richest and middle-income countries.
-This convergence was driven by improvements in the capital-output ratio (121%) and human capital (143%), which offset the divergence from TFP differences (-164%).
-The P50/P10 ratio increased by 42%, indicating divergence between middle-income and poor countries, primarily due to TFP differences (62%).
-B.5 Variance Decomposition
+Across all countries the estimated convergence rate lies between $-0.0062$ (output-side GDP, current sample) and $-0.0043$ (expenditure-side GDP, PSS sample); all coefficients are negative and statistically significant at the one-percent level.
+Restricting attention to economies outside SSA roughly triples the speed of convergence, with point estimates ranging from $-0.0155$ to $-0.0128$.
+It can thus be concluded that $\beta$-income convergence remains robust to the choice between RGDPO and RGDPE.
+Appendix C Additional Figures
+Figure 6: Decomposition Analysis (excluding SSA): P90/P50
+Notes: The figure decomposes the change in the log P90/P50 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for the periods 1980-2000 and 2000-2019 for countries outside Sub-Saharan Africa.
+Inputs (capital-output ratio and human capital) account for 66% of the convergence in the log P90/P50 ratio between 2000-2019 and for the entire convergence over 1980-2019.
+Figure 7: Decomposition Analysis (excluding SSA): P50/P10
+Notes: The figure decomposes the change in the log P50/P10 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for the periods 1980-2000 and 2000-2019 for countries outside Sub-Saharan Africa.
+Inputs (capital-output ratio and human capital) account for 32% of the convergence in the log P50/P10 ratio between 2000-2019 and for 70% of the convergence over 1980-2019.
+Figure 8: Decomposition Analysis (full sample): P90/P50
+Notes: The figure decomposes the change in the log P90/P50 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for the periods 1980-2000 and 2000-2019 using the full sample of countries.
+Inputs (capital-output ratio and human capital) account for 49% of the convergence in the log P90/P50 ratio between 2000-2019 and for the entire convergence over 1980-2019.
+Figure 9: Decomposition Analysis (full sample): P50/P10
+Notes: The figure decomposes the change in the log P50/P10 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for the periods 1980-2000 and 2000-2019 using the full sample of countries.
+During 2000-2019 inputs (capital-output ratio and human capital) diverged, whereas TFP exhibited a slight convergence, yielding a net divergence in the P50/P10 ratio.
+Over the entire horizon 1980-2019 TFP divergence explains 63 % of the overall income divergence, while input divergence accounts for a further 27 %.
+Figure 10: Decomposition Analysis (excluding SSA, $\alpha=1/3$): P90/P50
+Notes: The figure decomposes the change in the log P90/P50 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third.
+Inputs (capital–output ratio and human capital) account for 47% of the convergence in 2000–2019 and for the entire convergence over 1980–2000.
+Figure 11: Decomposition Analysis (excluding SSA, $\alpha=1/3$): P50/P10
+Notes: The figure decomposes the change in the log P50/P10 output ratio into the contributions of total factor productivity, the capital–output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third.
+Inputs account for 10% of the convergence in 2000–2019 and for 37% of the convergence over 1980–2000.
+Figure 12: Decomposition Analysis (full sample, $\alpha=1/3$): P90/P10
+Notes: The figure decomposes the change in the log P90/P10 output ratio into the contributions of total factor productivity, the capital–output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third.
+Capital convergence accounts for 28% of the convergence in 2000–2019.
+Over the full 1980–2019 horizon the ratio diverges overall, driven by large TFP divergence in 1980–2000; capital convergence mitigates but does not overturn this early divergence.
+Figure 13: Decomposition Analysis (full sample, $\alpha=1/3$): P90/P50
+Notes: The figure decomposes the change in the log P90/P50 output ratio into the contributions of total factor productivity, the capital–output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third.
+Inputs account for 42.6% of the convergence in 2000–2019 and for the entire convergence over 1980–2019.
+Figure 14: Decomposition Analysis (full sample, $\alpha=1/3$): P50/P10
+Notes: The figure decomposes the change in the log P50/P10 output ratio into the contributions of total factor productivity, the capital–output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third.
+In 2000–2019 overall output diverges despite TFP convergence because of divergence in capital inputs.
+Over 1980–2019 both TFP and capital diverge, reinforcing the long-run divergence in the P50/P10 ratio.
+C.1 Variance Decomposition
 This section complements the main accounting analysis by performing the accounting based on variance as a measure of income dispersion, which is more commonly used in the literature.
@@
 Furthermore, variance decomposition includes a term for the covariance between total factor productivity (TFP) and inputs, which can be difficult to interpret.
-For this exercise, I use a Cobb-Douglas production function, and a capital share parameter ($\alpha$) of 0.46, which is the observed mean in PWT10.01:
+For this exercise, I use a Cobb-Douglas production function, and a capital share parameter ($\alpha$) of 0.46, which is the observed mean in PWT10.01.
 Suppose we have a Cobb-Douglas production function.
@@
 This can be rewritten as
-$\frac{Y}{L}=A(\frac{K}{Y})^{\frac{\alpha}{1-\alpha}}(h)$
-$y=Ay^{kh}$
-where $y=\frac{Y}{L}$, $y^{kh}=\frac{K}{Y})^{\frac{\alpha}{1-\alpha}}(h)$. Consequently,
-$Var(\ln(y))=Var(\ln(A))+Var(\ln(y^{kh}))+2Cov(\ln(A),\ln(y^{kh}))$
-Table 9: Accounting for Variance Decline
-Years $\Delta_{t}Var(\ln(y))$ $\Delta_{t}Var(\ln(A))$ $\Delta_{t}Var(\ln(y^{kh}))$ $\Delta_{t}$ 2$\times$×Cov(A,$y^{kh}$)
-Outside SSA
-1980-2000 11 141 -126 86
-2000-2019 -39 27 36 36
-1980-2019 -27 -20 104 16
-Full Sample
-1980-2000 51 69 -20 51
-2000-2019 -19 66 43 -9
-1980-2019 31 71 -59 88
-Table 9 presents the variance decomposition of log income per worker ($\ln(y)$) for both countries outside Sub-Saharan Africa (SSA) and the full sample over three periods: 1980-2000, 2000-2019, and 1980-2019.
-For countries outside SSA, between 1980 and 2000, the variance of log income per worker increased by 11%, indicating a slight divergence in income levels among these countries.
-This increase was primarily driven by a substantial rise in the variance of log TFP, which surged by 141%, suggesting that differences in total factor productivity widened significantly during this period.
-The covariance term also increased by 86%, further contributing to divergence.
-However, the variance of log inputs decreased by 126%, indicating strong convergence in physical and human capital accumulation.
-Had it not been for the convergence in inputs, income variance would have been much higher, as the divergence from TFP differences and the covariance effect would have led to an even greater increase in income disparities.
-Between 2000 and 2019, the variance of log income per worker outside SSA decreased by 39%, reflecting a significant convergence in income levels.
-Decline in variance of inputs accounted for 36% of convergence.
-The contribution of inputs increases to 54% when including half the covariance term.
-The role of inputs is even more pronounced for the time period 1980-2019.
-During this period the variance of income decline by 27%, with decline in variance of inputs accounting for 104% of the decline, while TFP diverged during this period.
-This shows that inputs account for almost half of decline since 2000, and almost all of decline since 1980 in variance of income outside SSA.
-For the full sample of countries, between 1980 and 2000, the variance of log income per worker increased by 51%.
-This increase was primarily driven by a rise in the variance of log TFP by 69%.
-The covariance term also increased by 51%, further contributing to divergence.
-However, the variance of log inputs decreased by 20%.
-Between 2000 and 2019, the variance of log income per worker in the full sample decreased by 19%.
-Reduction in variance of TFP contributed to 66% of the decline, while reduction in variance of inputs contributed to 43% of decline.
-Covariance term increased by 9%, contributing to divergence.
-Over the entire period from 1980 to 2019, the variance of log income per worker in the full sample increased by 31%, indicating a net divergence in global income levels over the long term.
-This divergence was primarily driven by a substantial rise in the variance of log TFP, which increased by 71%, and an increase in the covariance term by 88%, both significantly contributing to divergence.
-However, the variance of log inputs decreased by 59%, indicating strong convergence in physical and human capital accumulation.
-B.6 Capital Accumulation in Sub-Saharan Africa
-It is puzzling to observe a persistent lack of convergence in Sub-Saharan African economies.
-Despite substantial improvements in policy outcomes across the region, as documented by [6], this has not translated into faster economic growth compared to economies outside Sub-Saharan Africa.
-[6] measures poor policy outcomes using indicators such as high inflation, significant black market premiums, and negative interest rates, among others.
-While these measures have improved considerably, growth in Sub-Saharan Africa remains sluggish relative to other regions.
-Although a comprehensive investigation into the reasons behind this lack of convergence is beyond the scope of the current study, I find that the slow rate of input accumulation is likely a key contributing factor.
-[17] further highlight capital flight from Sub-Saharan economies to the rest of the world, driven by economic and political uncertainties, which may exacerbate the region’s stagnation.
-Figure 2: Capital-Output Ratio across regions
-Figure 2 illustrates the mean capital-output ratio, weighted by population size, across different regions.
-While the capital-output ratio has continued to rise in regions outside Sub-Saharan Africa, Sub-Saharan Africa is the only region where a decline in the capital-output ratio is observed after 2000.
-This decline likely reflects the region’s challenges in accumulating capital, which may be linked to political instability, economic uncertainty, and capital flight.
-The capital-output ratio in Figure 2 is computed using a sample of countries for which investment data has been available since 1970 or earlier, ensuring consistent comparisons across regions.
-B.7 Influence of Venezuela
-Figure 3: Log GDP per capita (2017\$ PPP) for Venezuala
-Figure 3 illustrates the significant decline in Venezuela’s GDP per capita, measured in 2017 PPP dollars, following 2010.
-The sharp drop in Venezuela’s GDP per capita is a direct consequence of the country’s political and economic crisis.
-Given the unique circumstances surrounding Venezuela’s economic decline, it stands as an outlier in cross-country analyses.
-Therefore, I exclude Venezuela from the calculation of income variance, as variance is particularly sensitive to outliers.
+$\frac{Y}{L}=A\left(\frac{K}{Y}\right)^{\frac{\alpha}{1-\alpha}}h,$
+or
+$y=A\,y^{kh},$
+where $y=\frac{Y}{L}$ and $y^{kh}=\left(\frac{K}{Y}\right)^{\frac{\alpha}{1-\alpha}}h.$ Consequently,
+$Var(\ln(y))=Var(\ln(A))+Var(\ln(y^{kh}))+2\,Cov(\ln(A),\ln(y^{kh})).$
+Figure 15: Variance Decomposition Outside Sub-Saharan Africa
+Notes: The figure decomposes the change in the variance of log output per worker into the contributions of total factor productivity (TFP), the combined variance of physical- and human-capital inputs, and the covariance between TFP and inputs for the intervals 1980-2000 and 2000-2019 for countries outside Sub-Saharan Africa.
+Figure 16: Variance Decomposition, Full Country Sample
+Notes: The figure reports the same variance decomposition for the full set of countries.
+Figures 15 and 16 trace how the dispersion of income per worker has evolved by separating the variance of log output into three components: the variance of total factor productivity, the variance of combined physical- and human-capital inputs, and their covariance.
+For countries outside Sub-Saharan Africa, the variance of log output per worker rose by 11 percent between 1980 and 2000.
+TFP divergence contributed 141 percent of this increase, and the covariance between TFP and inputs added a further 86 percent, whereas convergence in physical and human capital offset 126 percent of the potential rise.
+In the subsequent period, 2000-2019, the variance fell by 39 percent: convergence in inputs accounted for 36 percent of that reduction, the fall in the covariance term accounted for another 36 percent, and TFP convergence accounted for the remainder.
+Taken together, these forces produced a net 27-percent decline in the variance over 1980-2019, with input convergence contributing 104 percent of the total reduction.
+For the full global sample, the variance of log output increased by 51 percent between 1980 and 2000, with TFP divergence responsible for 69 percent of the rise and the covariance term for 51 percent, while input convergence offset 20 percent.
+During 2000-2019 the variance contracted by 19 percent: TFP convergence contributed 66 percent of the decline, input convergence 43 percent, and a further expansion of the covariance term offset 9 percent.
+Over the entire horizon 1980-2019 the variance remained 31 percent above its initial level; TFP divergence accounted for 71 percent of this long-run increase and the covariance term for 88 percent, whereas input convergence reduced the total by 59 percent.
+Figure 17: Log GDP per capita (2017\$ PPP) for Venezuela
+Notes: The series displays a pronounced decline in Venezuela’s GDP per capita after 2010, reflecting the country’s political and economic crisis; Venezuela is treated as an outlier and excluded from the income-variance calculations.